### v1

# ZEUS: Zero-shot Embeddings for Unsupervised Separation of Tabular Data

Abstract
Clustering tabular data remains a significant open challenge in data analysis and machine learning. Unlike for image data, similarity between tabular records often varies across datasets, making the definition of clusters highly dataset-dependent. Furthermore, the absence of supervised signals complicates hyperparameter tuning in deep learning clustering methods, frequently resulting in unstable performance. To address these issues and reduce the need for per-dataset tuning, we adopt an emerging approach in deep learning: zero-shot learning. We propose ZEUS, a self-contained model capable of clustering new datasets without any additional training or fine-tuning. It operates by decomposing complex datasets into meaningful components that can then be clustered effectively. Thanks to pre-training on synthetic datasets generated from a latent-variable prior, it generalizes across various datasets without requiring user intervention. To the best of our knowledge, ZEUS is the first zero-shot method capable of generating embeddings for tabular data in a fully unsupervised manner. Experimental results demonstrate that it performs on par with or better than traditional clustering algorithms and recent deep learning-based methods, while being significantly faster and more user-friendly.
1 Introduction
Clustering remains a fundamental yet challenging task in unsupervised learning. It is particularly hard for tabular data, which inherently lacks the structured spatial or semantic properties of images or texts. Unlike for image clustering, where intrinsic visual similarities can guide cluster formation, defining meaningful similarities in tabular data is highly dataset-specific, complicating the generalization of clustering methods across diverse applications. Recent developments leveraging deep learning have demonstrated promise in generating richer representations for clustering tasks. However, these methods frequently suffer from instability due to their sensitivity to hyperparameter selection, a challenge exacerbated by the absence of supervised signals to guide optimization. Consequently, practitioners working with tabular data often resort to simpler, classical algorithms like k-means, despite their limited capacity for capturing complex underlying data structures, simply to avoid extensive manual tuning.
We address these challenges with ZEUS – a zero-shot transformer-based model for embedding new tabular datasets in a form convenient for unsupervised separation (=clustering), without the need for additional fine-tuning. Given a new dataset, ZEUS returns its transformed representation, where clusters can be easily discovered using simple methods, like k-means. Since it works as a zero-shot learner, it significantly reduces hyperparameter tuning complexity and the computation time, enabling effective clustering in seconds. It is a plug-and-play solution that requires no fine-tuning and runs in a single forward pass for new datasets.
Inspiration for ZEUS stems from the Prior-data Fitted Networks (PFNs) tabpfn , a recently introduced framework highlighting the potential of in-context learning for tabular data. PFNs operate by feeding a transformer contexts representing complete tasks, i.e., training data along with query samples for which the model predicts labels. Despite their impressive performance, they are limited to supervised problems. ZEUS extends this basic idea for unsupervised tasks by addressing the two key challenges of how to: (1) generate prior synthetic data with clear but non-trivial clustering structures for pre-training; and (2) encode prior clustering knowledge for new unlabeled datasets. Additionally, in contrast to TabPFN, ZEUS is a zero-shot model and during inference does not rely on any context labels.
Unlike traditional methods that optimize arbitrary heuristics (e.g., DEC), ZEUS approaches clustering by learning how to invert data generation processes. In particular, it is pre-trained on synthetic datasets to learn how to infer cluster assignments. The datasets have known latent structures, which enables supervised guidance for the training, and by generating diverse datasets we supply it with the prior knowledge about what can constitute a possible clustering structure, enabling effective generalization to new real-world datasets during inference. Figure 1 illustrates the key concepts of the method while Figure 2 presents representations generated by ZEUS.
Experiments demonstrate that ZEUS consistently matches or surpasses both classical clustering algorithms and recent deep learning-based approaches, establishing it as a powerful and practical tool for the unsupervised analysis of tabular data. Beyond its strong empirical performance, we also provide theoretical justification by showing that ZEUS fits the framework of Prior-data Fitted Networks, thereby reinforcing its theoretical soundness (see Section 2.3).
We supplement the paper with an appendix that includes background on prior-data fitted networks, details of the synthetic data generation process, and additional details of the experimental results. The code used in this paper is available at https://github.com/gmum/zeus
Figure 1: Schematic characterization of ZEUS: (left) synthetic datasets generation; (middle) pre-training on datasets with known labels; (right) deployment of a frozen model for real-world tasks.
2 Pre-training representation for clustering
In this section, we introduce our approach to zero-shot representation pre-training for clustering. First, we characterize the proposed probabilistic model and its training objective. Then, we explain the process of generating synthetic datasets for pre-training. Finally, we highlight the connection between our model, Bayesian learning and the framework for prior-data fitted networks.
Figure 2: t-SNE visualization for a sample synthetic dataset. The representation from ZEUS (right panel) significantly improves consistency with ground-truth classes and reveals a clearer data structure.
2.1 ZEUS method
Zero-shot pre-training.
Traditional machine learning models $f_{\theta}$ are trained by optimizing parameters $\theta$ to solve a specific fixed task, such as classification or clustering, over a fixed dataset. On the other hand, in-context learning (ICL) 10.5555/3600270.3602487 ; lee2023supervised ; han-etal-2023-understanding enables pre-trained models $f_{\theta^{\ast}}$ to adapt to new tasks or datasets without ever updating the pre-trained parameters $\theta^{*}$. Similarly, zero-shot learning allows models to address problems they were never explicitly trained on, without requiring any examples of the target task. Unlike ICL, which adapts through demonstration examples, zero-shot approaches rely entirely on knowledge encoded during pre-training. The pre-training of zero-shot models typically involves massive and diverse datasets that encourage the model to learn rich, generalizable representations of underlying patterns and relationships. This usually is done using transformers, as they both can process set-valued inputs (e.g., accept a whole dataset at once) and model complex dependencies.
Our approach involves pre-training a transformer $f_{\theta}$ on a diverse collection of datasets $\mathcal{D}$ such that $\mathcal{D}:=(\mathbf{x},\mathbf{y})$, where $\mathbf{y}=\{y_{i}\}$ denotes ground-truth labels (e.g., clusters) and $\mathbf{x}=\{x_{i}\}$ are feature vectors. A loss $\mathcal{L}$ encourages $f_{\theta}$ to assign accurate labels $\mathbf{y}$ to all data points in the target datasets $\mathcal{D}$:
$\theta^{\ast}=\arg\min_{\theta}\;\mathbb{E}_{p(\mathcal{D})}\bigl{[}\mathcal{L}\bigl{(}\mathbf{y},\,f_{\theta}(\mathbf{x})\bigr{)}\bigr{]}\approx\arg\min_{\theta}\;\frac{1}{N_{\mathcal{D}}}\sum_{\mathcal{D}\sim p(\mathcal{D})}\mathcal{L}\bigl{(}\mathbf{y},\,f_{\theta}(\mathbf{x})\bigr{)},$ (1)
where $p(\mathcal{D})$ is a process generating prior datasets which implicitly characterize a space of possible target tasks (here approximated with $N_{\mathcal{D}}$ Monte-Carlo samples). Note, the loss $\mathcal{L}$ may factorize over individual points as $\mathcal{L}(\mathbf{y},\hat{\mathbf{p}})=\sum_{i}\ell(y_{i},\hat{p}_{i})$, where $\hat{p}_{i}$ denotes a probabilistic prediction for an individual input $x_{i}$, but unlike for classification, for tasks such as clustering or anomaly detection, the predictions $\hat{p}$ must be made jointly for all inputs. Hence, $f_{\theta}(\mathbf{x})$ can not be decomposed into predictions for individual inputs $x_{i}$. Furthermore, in contrast to typical supervised ICL, training labels are not included in the context, as they are unavailable during inference. This makes the unsupervised learning problem significantly more challenging, since we cannot explicitly guide the model toward the desired structure in a given dataset.
Representation learning for probabilistic clustering.
We pre-train an encoder $f_{\theta}$ to map inputs $\mathbf{x}$ to representation vectors ${z}(x_{i})=f_{\theta}(\mathbf{x})_{i}\in\mathbb{R}^{D}=:S.$ We aim at positioning data points in $S$, so that they naturally form probabilistic clusters, and our goal is to make the encoder find better representations for target tasks. This objective we frame as maximizing the log-likelihood of correct cluster assignments. Since the ground-truth cluster assignments $y_{i}$ are available during pre-training, this is expressed as:
$\mathcal{L}_{prob}=-\sum_{i}\log p_{y_{i}}(x_{i}).$ (2)
For specifying $p_{y}(x)$ we draw an inspiration from Gaussian Mixture Models (GMMs) dempster1977maximum ; hansen1982large , where each cluster $k$ is represented by a Gaussian distribution centered at ${c}_{k}$, and data points are probabilistically assigned to clusters through probabilities $p_{k}(x_{i})$, based on their proximity to these centroids. Consequently, alongside the encoded representations, we need to define a set of $K$ cluster centroids (e.g., prototypical representations) $\{{c}_{k}\}_{k=1}^{K}$, which reside in the same representation space $S$. We then jointly optimize $\mathcal{L}$ w.r.t $\theta$ and $\{{c}_{k}\}$ such that the representations ${z}(x_{i})$ become structured in a way undisclosing underlying probabilistic structures in datasets.
We aim at higher probabilities $p$ for points closer these centroids. To achieve this, we first define scores $\alpha_{k}(x_{i})=-\|{z}(x_{i})-{c}_{k}\|^{2}$ to quantify the compatibility between representations and centroids ${c}_{k}$, and then transform them into cluster membership probabilities using softmax, yielding soft cluster assignments $p_{k}(x_{i})=\frac{\exp(\alpha_{k}(x_{i}))}{\sum_{j=1}^{K}\exp(\alpha_{j}(x_{i}))}.$ Note that the ground-truth cluster assignments $y_{i}$ are known during pre-training, and hence, the maximum likelihood centroids can be straightforward estimated as $\hat{c}_{k}=\frac{1}{N_{k}}\sum_{\{i:y_{i}=k\}}{z}(x_{i}),$ where $N_{k}$ denotes the number of data points assigned to the $k$-th cluster. Following this formulation, the optimization needs to performed only w.r.t the parameters $\theta$ of the neural network $f_{\theta}$, as the centroids are already specified by them.
Besides the centroids, we additionally consider another concept from GMMs, namely cluster priors ${\pi}_{k}$. Similar to centroids, the prior cluster frequencies $\hat{\pi}_{k}$ can be estimated from the training labels as the proportions of examples in the $k$-th cluster, and then incorporated directly into the probability calculation, resulting in the final form of our probabilistic scoring:
$p_{k}(x_{i})=\frac{\hat{\pi}_{k}\exp(\alpha_{k}(x_{i}))}{\sum_{j=1}^{K}\hat{\pi}_{j}\exp(\alpha_{j}(x_{i}))},\quad\textit{where}\quad\alpha_{k}(x_{i})=-\|{z}(x_{i})-\hat{c}_{k}\|^{2}.$ (3)
Remark 2.1.
ZEUS cluster assignments correspond to the membership probabilities in GMMs defined as: $r_{ik}=\frac{\pi_{k}\mathcal{N}({z}(x_{i})|{c}_{k},\Sigma_{k})}{\sum_{j=1}^{K}\pi_{j}\mathcal{N}({z}(x_{i})|{c}_{j},\Sigma_{j})},$ but with fixed diagonal covariance matrices $\Sigma_{k}:=I$, effectively forcing the encoder to learn circular clusters rather than elliptical ones.
Inference.
The structural relations learned during pre-training are leveraged during inference when the representations ${z}$ produced by the model $f_{\theta^{\ast}}$ are used to create a predictive distribution for new, previously unseen inputs $\mathbf{x}^{*}$ as $\hat{\mathbf{p}}^{*}:=p\bigl{(}\mathbf{y}^{*}\mid f_{\theta^{\ast}}(\mathbf{x}^{*})\bigr{)}.$ At this point, the labels $\mathbf{y}^{*}$ are unknown and therefore, cannot be used to estimate the centroiods nor the priors. Hence, to structure the obtained representation into clusters, we use a traditional learning algorithm, for example, similar to GMMs relying on Expectation-Maximization or simply k-means.
Regularization.
Although theoretically $\mathcal{L}_{prob}$ is sufficient to build a clustering representation, we experimentally verified that the introducing additional regularizers further improves the predictions in the inference (see Section 3.4).
First, we aim to ensure that the representations ${z}(x_{i})$ associated with a particular cluster are compactly distributed around their corresponding centroid to enhance intra-cluster cohesion. A point concentration regularizer that explicitly minimizes the distance between representations and the centroids, analogous to the k-means objective of minimizing within-cluster sum of squares, we define as
$\mathcal{L}_{cp}=\sum_{k}\sum_{i:y_{i}=k}\alpha_{k}(x_{i}).$ (4)
Second, to prevent the cluster centroids ${c}_{k}$ from collapsing towards similar points in the space, we add a centroid separation regularizer. We achieve this by maximizing the sum of squared distances between all distinct pairs of centroids. However, to avoid this term dominating the loss if centroids were pushed infinitely far apart, we cap the contribution of each pair’s squared distance at a predefined threshold $T$. The term to be minimized is thus:
$\mathcal{L}_{\textit{sep}}=-\sum_{k=1}^{K}\sum_{j=k+1}^{K}\min(\|\hat{c}_{k}-\hat{c}_{j}\|^{2},T).$ (5)
The final loss combines the main clustering objective with two regularization terms as
$\mathcal{L}=\mathcal{L}_{prob}+\lambda_{cp}\mathcal{L}_{cp}+\lambda_{sep}\mathcal{L}_{sep}$ (6)
where the hyperparameters $\lambda_{cp}\geq 0$ and $\lambda_{sep}\geq 0$ promote point concentration and control the relative importance of enforcing centroid separation, respectively. We used simply $\lambda_{cp}=\lambda_{sep}=1$.
2.2 Prior data for pre-training ZEUS
The key component of ZEUS is the data-generating prior $p(\mathcal{D})$. As it primarily affects the generalization ability of the pre-trained model $f_{\theta^{*}}$, it must be designed to cover diverse data distributions and various cluster configurations. In particular, we construct it as a latent variable model (LVM), assuming that each dataset $\mathcal{D}$ is sampled from a $K$-component mixture of distributions. Formally, the probability of a data point $x\in\mathcal{D}$ under this model is $p(x)=\sum_{k=1}^{K}p(y=k)\,p(x\mid y=k),$ where $p(y)$ is a categorical distribution, and $p(x\mid y=k)$ represents the $k$-th (continuous) component. Although the categories $y$ remain latent for real datasets, they are known during synthetic data generation and can serve as ground-truth labels – a property integral to pre-training ZEUS.
The number of categories $K$ we sample uniformly between 2 and 10, and the observations $x$ are sampled from multivariate Gaussian distributions. To control the complexity of datasets, we introduce a constraint that provides a sufficient separation between each pair of components. For each component, we generate between $50$ and $800$ samples. Since real data rarely consist of Gaussian clusters, we additionally transform the data points from each category using randomized ResNet-like neural networks to produce more realistic cluster shapes. We selected ResNets due to their properties, as they can define invertible transformations  (behrmann2019invertibleresidualnetworks, ), and this ensures that clustering structures will be preserved in outputs. Finally, we append a certain fraction of categorical features in one-hot encoding scheme to selected datasets, which is followed by an optional PCA reduction to keep the data dimension at the requested maximum level. Complete details of the prior-data generating process can be found in Appendix B.
The above strategy for generating clustering datasets illustrates the key idea behind ZEUS. Instead of clustering each dataset individually using an arbitrary loss criterion (as in $k$-means or DEC), ZEUS learns to perform clustering by inverting the data generation process. Although our synthetic datasets are limited by the selected family of distributions specified above, the proposed paradigm for zero-shot unsupervised learning for clustering is general.
2.3 Relation to Bayesian learning and Prior-Data Fitted Networks
The recently introduced framework of Bayesian inference through transformers demonstrates that neural networks pre-trained on synthetic datasets implicitly approximate Bayesian inference without explicitly computing posterior distributions bayes_in_context ; tabpfn . By framing the approach described in Section 2.1 as a Prior-Data Fitted Network (PFN), we show ZEUS implicitly performs approximate Bayesian averaging.
[…]
ZEUS instead of directly outputting $q_{\theta}(y|x,\mathcal{D}_{ctx})$, maps inputs $x$ to latent representation vectors ${z}(x)$. The probabilistic assignments $p_{y}(x)$ are then constructed from these vectors according to Eq. 3. This equation specifies a PPD for inference, but also defines a probability mass function for training:
Remark 2.2.
Eqs. 1, 2, and 3 constitute a valid Prior-Data Negative Log-Likelihood, equivalent to Eq. (2) from bayes_in_context with $\mathcal{D}_{ctx}=\emptyset$.
This follows by mapping $p_{y}(x):=q_{\theta}(y|x,\emptyset)$ and noting that Eq. 2 corresponds to the cross-entropy between the true labels $\{y\}$ and the probabilistic assignments $p_{y}(x)$. Intuitively, this formulation reinterprets probabilistic clustering with known labels as a classification task. Then, pre-training the transformer by minimizing the cross-entropy loss over prior-generated datasets remains identical to PFNs, and thus, we can conclude that ZEUS implicitly learns a Bayesian approximation through prior fitting.
Our method, however, deviates from traditional PFNs by imposing an explicit mixture-like structure on latent representations (Eq. 3), unlike more general PFNs:
Remark 2.3.
By enforcing the clustering structure, ZEUS may impose stronger assumptions on the PPD compared to vanilla PFNs. This structure might be suboptimal for classification tasks and the true PPD may not belong to the family of attainable solutions.
As explained above our clustering-based extension is theoretically sound, nonetheless, the assumption could limit representational expressivity compared to the more flexible transformers employed for the original PFNs tabpfn . On the other hand, the enforced structures shall be more appropriate for unsupervised tasks.
3 Experiments
This section provides an experimental evaluation of the clustering performance of our method.
3.1 Experimental setup
Model architecture: ZEUS relies on a transformer architecture similar to TabPFN (tabpfn, ). It consists of 12 attention blocks, each with 6 heads and a token dimension of 512, employing GeLU as the activation function. Following the TabPFN approach, each data point is linearly embedded prior to the transformer component and then treated as a token.
Pre-training: In the pre-training phase, we sample datasets from the mixture of Gaussians and transformed mixtures in equal proportions (1:1), generating 1000 unique dataset batch samples for each epoch. For training, we employ the Adam optimizer along with a cosine learning rate scheduler with warm-up, using a learning rate of 2e-5. The plot illustrating model improvements during the pre-training process is available in Appendix C.
Inference: During inference, preprocessing of each dataset involves standardizing numerical features, followed by scaling them to the range $[-1,1]$, whereas categorical features are transformed using one-hot encoding. The input size of our model is fixed to 30. For datasets with lower dimensionality, we pad the missing positions with zeros, while for higher-dimensional datasets, we reduce the number of input features via Principal Component Analysis (PCA). Unless stated otherwise, at inference we use k-means applied to the normalized (=scaled to $[-1,1]$) transformer output in order to obtain clusters from our learned representation.
Datasets: For evaluation, we consider three groups of datasets: real datasets from OpenML (Real), synthetic mixtures of Gaussians (Syn. Gauss.), and synthetic mixtures of Gaussians transformed by ResNet-like neural networks (Syn. Transf.). Both types of synthetic datasets are augmented with categorical variables. The process of generating synthetic datasets is described in Section 2.2.
Each dataset contains at most 2000 samples as per model design and due to memory limitations. The study covers 34 real datasets, selected based on their clustering feasibility, defined as ARI $\geq 0.4$ achieved by at least one of the methods. Additionally, 20 synthetic datasets of each type were generated from the same prior as in the pre-training phase, but with a different random seed to ensure a fair comparison. Detailed statistics of the datasets, such as the number of numerical and categorical features, are provided in Appendix D.
Baselines: We compare ZEUS against a wide spectrum of state-of-the-art clustering methods used for tabular data. It includes $k$-means (KM), Gaussian Mixture Model (GMM), and deep-learning methods based on autoencoder (AE) architectures, including DEC xie2016unsuperviseddeepembeddingclustering , IDEC guo_improved_2017 , and G-CEALS rabbani2025deep . We additionally consider also $k$-means and GMM applied in the autoencoder latent space (respectively referred to as AE-KM and AE-GMM) to compare the clustering quality of our representation against the one obtained from the autoencoder.
For AE-based baselines, we employ the standard configuration used in the prior literature, i.e., an architecture comprising of hidden layers with sizes [500, 500, 2000]. We used the latent dimension of 20, which we verified experimentally as the best value. Further details on the hyperparameters and code repositories for the baselines are provided in Appendix E.
Reporting: In the main text, we report only the results aggregated for each group of datasets (Real, Syn. Gauss, Syn. Transf.), while detailed results for individual datasets can be found in Appendix F. For the reader’s convenience, we bold the best results and underline the second-best ones.
3.2 How effective is ZEUS for clustering?
To evaluate performance of the clustering methods, we employ a standard evaluation procedure, where clusters identified by models are expected to correspond to undisclosed ground-truth classes. We use the Adjusted Rand Index (ARI) for quantitative evaluation, and to improve readability, we scale the ARI values by a factor of 100, where 100 indicates perfect clustering and values near 0 represent random grouping. Table 1 presents a summary of ARI scores averaged over 5 random seeds and all datasets within each dataset group. Additionally, Table 2 displays the average rankings for each of the methods.
Table 1: Clustering quality (ARI) of ZEUS versus competing methods (higher is better).
KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
Real 55.54 48.49 51.43 53.56 55.93 54.57 40.37 57.43
Syn. Gauss. 89.90 76.93 81.26 81.40 89.36 82.57 62.84 89.03
Syn. Transf. 75.04 75.88 60.45 71.29 79.94 61.26 49.17 86.33
Table 2: Average rank of the methods used in the benchmark (lower is better).
KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
Real 3.90 4.57 4.72 4.29 3.96 4.31 6.65 3.60
Syn. Gauss. 2.60 3.35 5.40 5.45 3.10 5.50 7.70 2.90
Syn. Transf. 4.45 3.20 6.20 4.25 3.00 5.70 7.05 2.15
We observe that ZEUS achieves competitive performance for all three groups of datasets. In particular, it achieves the best average ARI and the best rank for both the OpenML and Synthetic Transformed datasets, e.g., for the most challenging clustering scenarios. Notably, ZEUS outperforms the second-best method, DEC, by more than 6 percentage points, and the classical baselines by over 10 percentage points on the Synthetic Transformed datasets. The results on the Real datasets further demonstrate that ZEUS effectively generalizes the knowledge acquired during the pre-training stage on synthetic data to general data distributions. For the simplest Gaussian datasets, ZEUS ends up in top-3, with results only slightly below those of k-means and DEC. In terms of average rank it gets the second place.
Figure 3: Clustering time vs. input size.
Among the AE-based methods, DEC performs the best, while G-CEALS consistently demonstrates the weakest performance across all datasets. K-means and GMM, as representatives of the classical methods, perform reasonably well. In particular, k-means consistently lands among the best performing methods, which justifies its broad adoption among practitioners, despite being one of the most basic approaches. The detailed results including scores for individuals datasets corresponding to the averages in Tables 1 and 2 can be found in Appendix F.1
Finally, Figure 3 illustrates how the examined clustering methods scale with the increasing number of input data points. ZEUS maintains almost constant time while being only slightly slower than the basic k-means. It shows that the overhead from creating the representations by ZEUS is minimal. On the other hand, the remaining deep clustering algorithms require significantly more time and scale poorly with the increasing input size.
3.3 Are ZEUS’s predictions calibrated?
For applications where uncertainty quantification is as crucial as the clustering decisions themselves, assessing calibration is particularly relevant. Having demonstrated the strong performance of ZEUS for hard clustering, we now examine whether its probabilistic foundations yield well-calibrated soft assignments. We compare it against competing approaches using the Brier score, which measures the accuracy of predicted probabilities. Unlike the previously used ARI, it penalizes both incorrect cluster assignments and poorly calibrated confidence scores, thereby providing a more comprehensive evaluation.
The Brier score is a supervised metric, meaning its direct computation for unsupervised clustering tasks is not straightforward. However, when ground-truth classes are available and the number of clusters matches the number of classes, a one-to-one mapping between clusters and classes can be established using the Hungarian algorithm (kuhn1955hungarian, ; kuhn1956variants, ), which aims to maximize the total agreement between cluster-class pairs. The cost matrix $A$ for this assignment problem is $A_{jc}=\sum_{i=1}^{N}p_{ij}\cdot Y_{ic}$, where $Y_{ic}=1$ if data point $i$ belongs to class $c$, and $Y_{ic}=0$ otherwise.
Table 3 reports the average Brier score computed over 5 seeds for all datasets in respective groups. For the analysis, we used the vanilla variant of ZEUS without regularization, which is paired with the GMM clustering since k-means does not provide soft assignments. The covariances were constrained to be identities as implied by Eq. 3. All competing baselines, except for k-means, provide probabilistic cluster assignments, making the Brier score calculation straightforward for them. For k-means, we used one-hot encoding to represent its assignments as probabilities.
Table 3: Soft clustering quality of ZEUS vs. baselines, measured by Brier score (lower is better).
KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
Real 0.4366 0.4799 0.4643 0.4679 0.3941 0.3671 0.4722 0.3817
Syn. Gauss. 0.0970 0.3110 0.2073 0.2484 0.3943 0.2308 0.3946 0.1269
Syn. Transf. 0.2803 0.2566 0.4535 0.3140 0.4638 0.3892 0.4951 0.1796
ZEUS demonstrates outstanding performance, achieving the best results for the Synthetic Transformed datasets while ranking second on both the OpenML and Synthetic Gaussian collections. The benefits of using ZEUS representations are especially visible in comparison against the vanilla GMM. Although the GMM achieves the second-best score on transformed data, its performance across the remaining datasets is merely modest. Among the remaining baselines, IDEC exhibits an interesting pattern: despite relatively weak clustering performance in Table 1, it achieves the top position on real datasets according to the Brier score and shows marked improvement on synthetic data. DEC displays the opposite tendency, with calibration results significantly inferior to its strong ARI performance. While G-CEALS remains generally weaker than other methods, its calibration performance consistently exceeds its clustering results presented in Table 1. The K-means algorithm, both in its standard implementation and when applied to autoencoder embeddings, yields impressive scores on Gaussian-categorical mixture datasets – likely due to the inherent separability of these data structures, which allows even simple one-hot probability estimates to produce well-calibrated predictions.
3.4 How helpful is regularization for ZEUS?
Representations obtained by minimizing the basic loss $\mathcal{L}_{prob}$ can be further improved by regularizing the optimization process to structure the latent representations. In Table 4, we examine how different combinations of the $\mathcal{L}_{prob}$ loss with the regularizers $\mathcal{L}_{cp}$ and $\mathcal{L}_{sep}$ affect the final performance across various datasets. To ensure a fair comparison, all models were evaluated using a data generator with fixed (identical for all) settings.
Table 4: Regularisation impact, measured by ARI (higher is better).
$\mathcal{L}_{prob}$ $\mathcal{L}_{prob}+\mathcal{L}_{sep}$ $\mathcal{L}_{prob}+\mathcal{L}_{cp}$ $\mathcal{L}_{prob}+\mathcal{L}_{sep}+\mathcal{L}_{cp}$
Real 44.80 51.60 48.65 57.43
Syn. Gauss. 83.37 81.88 90.59 89.03
Syn. Transf. 79.85 79.29 88.58 86.33
Our main model, which during pre-training incorporates both regularizers, exhibits a clear performance advantage on real datasets and consistently holds second place for the synthetic ones. Table 4 shows that a model pre-trained using only $\mathcal{L}_{prob}$ and its variant with $\mathcal{L}_{sep}$ included are insufficient for properly separating the transformer’s representations for the clustering task, as they both struggle with clustering the synthetic datasets encountered during pre-training. On the other hand, the results for $\mathcal{L}_{prob}+\mathcal{L}_{cp}$ imply a positive impact by the compact loss component, $\mathcal{L}_{cp}$, especially for synthetic data, which however does not translate to strong performance on the OpenML benchmark. We conclude that all three loss components are necessary for good performance, with $\mathcal{L}_{cp}$ being crucial for pre-training and $\mathcal{L}_{sep}$ important for real data.
This mismatch between performance on real and synthetic datasets suggests a potential prior misspecification, which is then mitigated by the regularizers. In particular, among the real datasets, there may be some that are not well explained by the data generation process used during pre-training. Hence, for future work, one may want to explore alternative data-generating priors that may be more appropriate for these outliers.
4 Related work
Tabular data clustering.
Traditional clustering algorithms, like k-means macqueen1967some , GMM dempster1977maximum , or hierarchical methods murtagh2012algorithms , have widespread applications across data mining, bioinformatics, customer segmentation, and anomaly detection. However, these methods often rely on predefined distance metrics and fail to capture complex, non-linear relationships, making them suboptimal for high-dimensional and heterogeneous tabular datasets.
A pioneering Deep Embedded Clustering (DEC) xie2016unsuperviseddeepembeddingclustering improves the target data representation by training the autoencoder and computing soft assignments in its latent space via Student’s $t$-distribution. In a concurrent work yang2017towards , the authors perform joint dimensionality reduction using AE and k-means clustering in the latent space of AE. Improved DEC (IDEC) guo_improved_2017 extends DEC by jointly optimizing reconstruction and clustering objectives, while spectral variations replace k-means steps with graph-based updates 8970928 . G-CEALS rabbani2025deep replaces the $t$–distribution assumption with multivariate Gaussian clusters. To eliminate the assumption of explicit distribution prior, DEPICT dizaji2017deepclusteringjointconvolutional attaches a softmax layer on top of an embedding network and train with cross-entropy loss. Finally, IDC svirsky2023interpretable predicts interpretable cluster assignments at the instance and cluster levels.
However, most of these deep learning models require careful hyperparameter tuning and early stopping which is unrealistic in the fully unsupervised setting due to the lack of labels 9506051 ; sadeghi2024deepclusteringselfsupervisionusing . Moreover, the optimization process has to be performed on each dataset, which is often time-consuming. Although multiple deep clustering approaches are currently in use, most of them are designed for texts or images znalezniak2023contrastive ; cai2023semantic ; li2024image and cannot be directly adapted to tabular data due to the lack of a dominant neural architecture for heterogeneous tabular inputs grinsztajn2022tree ; mcelfresh2023neural .
Representation learning for tabular data.
Self-supervised learning (SSL) has been transformative in domains like vision and language chen2020simple ; grill2020bootstrap but has struggled to show similar success in tabular data yoon2020vime ; zhu2023xtab . Large diversity of data and lack of pre-defined correlation between features prevents from designing universal pretext tasks or augmentations as well as transferability between domains zhao2024comparison . In-context learning seems to be a promising direction in representation learning, which enables a single pre-trained model to adapt to new tasks in a zero-shot setting brown2020languagemodelsfewshotlearners . TabPFN treats small tabular datasets as contexts of (features, labels) and achieves state-of-the-art classification in one forward pass tabpfn . Although theoretical analyses reveal that transformers can implicitly implement algorithms such as gradient descent in context, their use is currently restricted to supervised problems garg2023transformerslearnincontextcase .
5 Conclusion
In this paper, we presented ZEUS, a zero-shot transformer-based model that enables effective and efficient clustering of tabular data without the need for fine-tuning or extensive hyperparameter search. By pre-training on synthetic datasets with known latent structures, ZEUS learns generalizable representations that help simple clustering algorithms to uncover meaningful structures. Our experiments show that ZEUS consistently matches or outperforms both classical and deep learning-based clustering methods, offering a practical and theoretically grounded solution for unsupervised analysis of tabular data.
Limitations.
Since ZEUS is technically based on the TabPFN architecture, it inherits some of its drawbacks: a maximum number of input features and samples. However, the recently introduced TabPFN v2 (hollmann2025accurate, ) showed that tabular transformers can process larger datasets with a negligible increase in computational time. Moreover, our experimental results demonstrate that even if the dimension of input data exceeds the fixed value of 30, applying PCA does not significantly hurt the clustering performance.
The final performance of ZEUS heavily relies on the synthetic data used in pre-training. While we release a basic version of ZEUS, which encodes certain assumptions about data clusters, one can adjust the pre-training stage using different datasets. Finally, ZEUS does not cluster data itself, but constructs a convenient embedding space in which clustering can be performed using basic algorithms.
Acknowledgments and Disclosure of Funding
This research is part of the project No. 2022/45/P/ST6/02969 co-funded by the National Science Centre and the European Union Framework Programme for Research and Innovation Horizon 2020 under the Marie Skłodowska-Curie grant agreement No. 945339.
For the purpose of Open Access, the authors have applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
[Uncaptioned image] [Uncaptioned image]
The research of M. Śmieja and P. Marszałek was supported by the National Science Centre (Poland), grant no. 2023/50/E/ST6/00169.
The research of J. Tabor was supported by the National Science Centre (Poland), grant no. 2023/49/B/ST6/01137.
Witold Wydmański is supported by the Ministry of Science grant no. PN/01/0195/2022 and NCN Sonata BIS grant number 2020/38/E/NZ2/00598.
Appendix A Background on Prior-Data Fitted Networks
The recently introduced framework of Bayesian inference through transformers demonstrates that neural networks pre-trained on synthetic datasets implicitly approximate Bayesian inference without explicitly computing posterior distributions [23, 15]. The pre-trained transformers are known as Prior-Data Fitted Networks (PFNs). The pre-training involves synthetic prior fitting, wherein a transformer network is trained offline on numerous datasets generated from a predefined prior distribution over tasks. This pre-training procedure allows the transformer to implicitly encode a Bayesian posterior predictive distribution by optimizing a cross-entropy loss between its predictions and the synthetic data labels.
Formally, pre-training is carried out as follows: given a prior distribution $p(\mathcal{D})$ over (synthetic) datasets $\mathcal{D}$, a PFN network parameterized by $\theta$ is trained to minimize the negative log-likelihood of predicting held-out labels within datasets sampled from this prior. The associated loss function, known as the Prior-Data Negative Log-Likelihood (Prior-Data NLL), is explicitly defined as:
$\mathcal{L}_{\textit{PFN}}(\theta)=\mathbb{E}_{\mathcal{D}_{ctx}\cup\{(x,y)\}\sim p(\mathcal{D})}[-\log q_{\theta}(y|x,\mathcal{D}_{ctx})],$
where each dataset $D_{ctx}$ and data point $(x,y)$ are sampled from the predefined prior distribution $p(D)$. As shown formally in [23], minimizing this loss is mathematically equivalent to minimizing the expected cross-entropy between the predictive distribution $q_{\theta}(y|x,D_{ctx})$ and the true posterior predictive distribution (PPD) derived from the prior. Specifically, minimizing the Prior-Data NLL is equivalent to minimizing the expected Kullback–Leibler divergence between the network’s predictive distribution and the true PPD $p(y|x,D_{ctx})$:
$\mathcal{L}_{\textit{PFN}}(\theta)=\mathbb{E}_{\mathcal{D}_{ctx},x\sim p(D)}[H(p(\cdot|x,\mathcal{D}),q_{\theta}(\cdot|x,D_{ctx}))],$
where $H$ denotes the cross-entropy. Thus, optimality of the predictive distribution $q_{\theta}$ implies matching it exactly to the true Bayesian posterior predictive distribution, provided that the parametric family of distributions defined by the transformer is sufficiently expressive [23].
Within this in-context learning as Bayesian inference framework, transformers approximate Bayesian averaging implicitly. Given a training dataset $\mathcal{D}_{ctx}=\{(x_{i},y_{i})\}_{i=1}^{n}$, a new input $x$, and a prior over hypotheses $\phi$, the Bayesian posterior predictive distribution is formally expressed as:
[…]
integrating over all hypotheses $\phi\in\Phi$. PFNs approximate this integral implicitly through pre-training on prior-generated datasets, yielding a distribution $q_{\theta}(y|x,\mathcal{D}_{ctx})$ directly from forward propagation of the network conditioned on dataset $\mathcal{D}_{ctx}$ [23, 15].
Appendix B Details of synthetic data generation process
We use three types of probabilistic models to generate data from a category:
  1. 1.
    Gaussian. The continuous features of each cluster follow a multivariate Gaussian distribution with carefully constructed means and covariance matrices. Cluster means are placed with a minimum Wasserstein-2 distance constraint to ensure adequate separation between clusters and to prevent trivial overlap cases. The exact minimum distance value varies between 0.5 and 1.0 and is randomly selected for each component independently in order to promote data diversity. The covariance structure for each Gaussian is generated through eigendecomposition, with eigenvalues sampled from a predefined range of $[0.005,0.05]$ to control the shape and orientation of the clusters. Additionally, to guarantee the presence of both full-rank Gaussians in higher dimensions and degenerate ones in lower dimensions, extra conditions are introduced to narrow the range of eigenvalues. These constraints are activated with probabilities of 0.25 and 0.2, respectively. The process of constructing Gaussian mixture is incremental. Starting with a mixture containing $k$ components, the addition of a new cluster involves initially placing it at position 0 and then shifting it in a randomly chosen direction until the distance requirement is satisfied. This entire approach allows us to model a variety of cluster geometries, from spherical to highly elongated elliptical shapes.
  2. 2.
    Categorical. We also incorporate categorical features alongside continuous ones by sampling from categorical distributions that are biased toward certain categories for each cluster. To generate these varied categorical probability patterns, we use the Dirichlet distribution. The resulting categorical variables are then converted to one-hot encoding and combined with the continuous features, producing mixed-type datasets that more closely resemble real-world tabular data. The probability of including categorical features is controlled by a $categorical\_chance$ parameter, set to 0.3. Up to 3 categorical variables may be added, each having between 2 and $max\_categories$ possible values, defined as 5 .
  3. 3.
    NN transformed. To create more complex, non-linearly separable cluster structures, we apply transformations to the numerical features using random neural networks with 3 to 6 layers. These transformations map the original data through several non-linear operations while preserving cluster identity information, producing datasets with more challenging decision boundaries. Following the approach of Invertible Residual Networks [1], we constrain the spectral norm of each transformation layer to be less than 1. In addition, standardization is applied between residual layers to ensure more stable transformations. To help preserve cluster separability, we append one-hot vectors encoding the component identity of each data point to the numerical variables before passing them through the random neural network. After the transformation, these extra dimensions are removed using PCA, restoring the data to its original dimensionality. This comprehensive approach enables the simulation of intricate, non-linear data manifold structures often presented in real-world clustering problems without introducing degenerate configurations.
Afterwards, all features undergo standardization and scaling to ensure numerical stability during training. Continuous features are normalized to the range $[-1,1]$, while ensuring that the relative separation between clusters is preserved.
By training on this diverse collection of synthetic datasets-ranging from well-separated Gaussian clusters to complex, transformed manifolds with mixed feature types-our model learns to identify meaningful cluster structures across a wide spectrum of data distributions. This comprehensive prior fitting enables ZEUS to adapt to previously unseen datasets at inference time without additional training.
Appendix C Pre-training plots
Figure 4: Visualization of pre-training process
Figure 4 presents two plots of average ARI over 200 synthetic validation datasets throughout 300 pre-training epochs. The first plot corresponds to the Gaussian-categorical datasets referred to as Synthetic Gaussian, while the second illustrates their NN transformed variants, named Synthetic Transformed. The data generation procedures are described in more detail in Appendix B. Both plots illustrate that the quality of the ZEUS representation improves over time, indicating that the model is learning new patterns, as evidenced by the increasing ARI during the pre-training process.
Appendix D Statistics of datasets used in experimental study
Table 5: Real/OpenML datasets statistics.
ID # Instances # Numerical # Categorical Dimension # Categories # Classes
features features (one-hots)
14 2000 76 0 76 0 10
15 699 9 0 9 0 2
16 2000 64 0 64 0 10
18 2000 6 0 6 0 10
22 2000 47 0 47 0 10
35 366 1 33 130 129 6
51 294 6 7 25 19 2
53 270 13 0 13 0 2
56 435 0 16 32 32 2
61 150 4 0 4 0 3
187 178 13 0 13 0 3
377 600 60 0 60 0 6
458 841 70 0 70 0 4
481 209 7 1 14 7 2
694 310 8 0 8 0 9
721 200 10 0 10 0 2
733 209 6 0 6 0 2
745 159 14 1 20 6 2
756 159 15 0 15 0 2
796 209 6 1 36 30 2
820 235 12 0 12 0 2
840 205 17 8 68 51 2
854 158 5 2 14 9 2
1462 1372 4 0 4 0 2
1495 250 0 6 18 18 2
1499 210 7 0 7 0 3
1510 569 30 0 30 0 2
1523 310 6 0 6 0 3
4153 180 66 0 66 0 6
40496 500 7 0 7 0 10
40682 215 5 0 5 0 3
40705 959 42 2 44 2 2
42261 150 4 0 4 0 3
42585 344 4 2 10 6 3
Table 6: Synthetic gaussian datasets statistics.
ID # Instances # Numerical # Categorical Dimension # Categories # Classes
features features (one-hots)
0 1337 16 0 16 0 8
1 1383 23 0 23 0 8
2 1421 8 2 14 6 7
3 992 9 1 12 3 9
4 1314 8 0 8 0 9
5 1497 16 2 23 7 8
6 1370 2 0 2 0 6
7 1646 4 3 16 12 8
8 1520 11 0 11 0 8
9 1537 18 0 18 0 5
10 825 26 0 26 0 2
11 1112 9 0 9 0 5
12 1093 15 0 15 0 8
13 742 6 0 6 0 3
14 1595 11 2 18 7 7
15 1417 14 0 14 0 6
16 1787 28 0 28 0 5
17 764 19 0 19 0 4
18 889 25 0 25 0 5
19 1660 28 0 28 0 9
Table 7: Synthetic transformed datasets statistics.
ID # Instances # Numerical # Categorical Dimension # Categories # Classes
features features (one-hots)
0 1337 16 0 16 0 8
1 1627 30 0 30 0 9
2 1631 11 0 11 0 7
3 1891 15 1 17 2 10
4 1142 3 0 3 0 4
5 1222 24 0 24 0 9
6 953 6 0 6 0 6
7 1508 9 0 9 0 10
8 840 7 3 15 8 5
9 1745 14 0 14 0 9
10 1618 23 0 23 0 6
11 1432 13 0 13 0 9
12 1860 9 0 9 0 9
13 563 10 0 10 0 2
14 1033 6 2 13 7 3
15 750 2 0 2 0 4
16 1451 14 2 20 6 10
17 679 30 0 30 0 2
18 859 22 1 25 3 2
19 1493 11 0 11 0 7
Tables 5, 6, and 7 provide detailed information about the datasets used in the experimental analysis. Each table includes the number of set instances, the number of categorical and numerical features, the total dimensionality, and the overall number of categories, represented by the length of the one-hot encoded vectors. Additionally, the tables report the number of classes that each dataset contains. In Table 5, the ID column corresponds to the OpenML ID, while in the remaining tables it functions solely as an index.
Appendix E Baselines
The evaluation of baseline models is based on the following libraries and GitHub repositories:
  1. 1.
    scikit-learn - used for k-means and GMM
  2. 2.
    https://github.com/vlukiyanov/pt-dec - implementation of the DEC
  3. 3.
    https://github.com/dawnranger/IDEC-pytorch - source code for the IDEC
  4. 4.
To ensure fair comparison, hyperparameters are chosen to maximize the performance of each method with respect to their overall average rank across 5 random seeds. For this reason, all numerical features are preprocessed using a standard scaler prior to the training phase.
Most parameters of the k-means and GMM methods are left at their default values. Only the $n\_init$ option was increased to 100 for k-means and 50 for GMM in order to improve stability.
As mentioned in Section 3.1, a standard autoencoder with hidden layers $[500,500,2000]$ is used for each AE-based method. The resulting network is first pre-trained for 1000 epochs and then fine-tuned for up to 1000 additional epochs, separately for each dataset and model. Following the GCEALS evaluation procedure, multiple latent dimension sizes $[5,10,15,20]$ were tested across all considered methods. The results indicate that the default latent dimension of 10 does not yield the best performance; instead, a dimension of 20 generally performs better. Changing other hyperparameters, including the learning rate, optimizer, and clustering loss weight $\gamma$, generally did not lead to improved results. Therefore, the remaining parameters were left at their default values as proposed by the authors of the respective repositories.
Table 8: Real/OpenML datasets - ARI
ID KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
14 38.74 45.90 40.94 45.36 49.60 36.78 45.04 50.56
15 82.85 71.38 61.44 71.74 86.29 87.48 24.38 81.28
16 55.62 64.75 61.66 69.97 68.56 55.61 55.27 74.03
18 54.86 46.88 51.45 51.47 50.00 53.54 49.93 51.63
22 35.48 47.39 31.09 50.80 50.02 35.69 28.38 56.05
35 70.68 70.68 70.22 70.48 68.63 64.58 56.57 85.12
51 28.35 -2.54 37.82 34.13 38.13 34.58 35.13 43.66
53 45.21 5.17 39.14 40.62 43.50 41.91 21.32 35.76
56 57.79 58.49 57.10 58.31 56.97 55.20 55.66 66.41
61 62.01 90.39 59.79 60.21 57.84 60.14 51.37 85.15
187 89.75 93.09 82.87 84.86 85.02 85.10 55.67 88.19
377 56.70 58.47 62.06 59.52 63.03 59.01 59.79 55.30
458 95.08 96.97 64.33 75.70 94.58 63.29 68.31 99.19
481 58.51 2.77 36.84 48.03 57.28 52.93 19.67 8.68
694 35.91 43.30 35.26 36.56 26.30 36.02 30.76 33.38
721 43.28 43.28 13.12 25.12 17.75 28.77 -0.03 43.28
733 50.97 46.29 38.39 55.11 67.47 73.51 26.26 74.97
745 57.63 55.44 78.35 75.52 46.37 50.94 57.81 73.85
756 55.70 4.46 72.81 53.40 58.60 65.58 44.14 50.31
796 49.62 2.26 76.54 52.19 50.07 55.92 43.40 13.99
820 50.16 36.77 44.29 49.24 51.21 49.30 24.51 28.87
840 39.05 48.37 53.60 46.72 49.14 39.16 27.95 27.66
854 76.14 -0.21 62.33 60.83 35.05 70.52 72.25 76.14
1462 1.32 0.31 10.03 10.17 1.00 3.80 8.48 92.28
1495 96.81 98.40 72.74 82.80 96.81 91.28 77.51 7.35
1499 77.33 62.99 73.68 75.28 69.99 78.48 29.60 82.40
1510 67.07 78.02 63.73 59.77 71.32 69.84 69.47 74.26
1523 21.16 41.21 25.48 24.57 28.16 23.05 18.50 19.60
4153 55.78 58.03 59.15 46.23 56.41 58.44 39.43 62.29
40496 54.02 34.79 42.35 37.51 49.74 42.35 37.33 32.25
40682 58.32 86.29 57.89 59.22 57.39 86.05 40.58 53.15
40705 43.49 37.92 0.73 30.63 47.90 46.28 10.71 45.45
42261 62.01 90.39 60.13 58.15 59.78 56.78 49.09 85.15
42585 60.84 30.52 51.26 60.71 91.79 43.40 38.30 95.05
Mean 55.54 48.49 51.43 53.56 55.93 54.57 40.37 57.43
Mean-rank 3.90 4.57 4.72 4.29 3.96 4.31 6.65 3.60
Top-3 16 14 9 8 18 13 2 21
Top-1 4 8 4 0 3 1 0 12
Table 9: Synthetic gaussian datasets - ARI
ID KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
0 86.16 82.13 83.44 79.66 87.06 77.26 58.57 87.65
1 88.66 87.36 81.13 79.61 85.84 75.62 54.10 72.04
2 81.30 26.46 71.57 78.04 86.65 78.91 48.59 82.96
3 95.02 36.77 92.03 76.44 92.74 89.81 62.67 95.76
4 90.64 92.97 82.65 77.80 91.61 84.51 69.70 92.85
5 89.83 25.85 86.52 69.06 90.01 88.31 67.70 90.72
6 96.21 98.07 85.99 94.26 94.24 92.39 73.15 97.89
7 92.59 30.52 81.73 79.06 91.35 89.80 58.10 93.60
8 88.29 89.62 81.15 76.53 87.56 78.37 50.30 84.00
9 93.36 98.54 68.58 85.42 91.14 85.29 56.89 94.55
10 79.26 93.32 72.75 90.90 72.54 73.51 48.47 87.32
11 93.60 98.48 90.77 89.11 95.03 92.19 64.24 97.45
12 86.82 77.45 73.64 71.40 86.04 77.93 58.45 86.04
13 96.77 98.07 94.05 90.15 96.11 85.66 78.20 96.17
14 90.73 43.20 79.65 69.09 92.68 77.93 67.84 94.60
15 93.58 96.23 85.74 89.30 94.43 89.41 72.88 96.68
16 94.70 92.45 90.98 89.41 94.53 88.14 85.86 87.19
17 85.26 87.10 62.23 79.52 84.08 75.00 42.55 86.29
18 88.77 94.83 84.47 84.74 87.39 83.46 67.96 83.01
19 86.54 89.11 76.18 78.46 86.06 67.99 70.50 73.85
Mean 89.90 76.93 81.26 81.40 89.36 82.57 62.84 89.03
Mean-Rank 2.60 3.35 5.40 5.45 3.10 5.50 7.70 2.90
Top-3 16 13 0 1 15 0 0 15
Top-1 3 10 0 0 1 0 0 6
Table 10: Synthetic transformed datasets - ARI
ID KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
0 60.56 79.33 46.31 68.98 76.93 46.35 29.42 87.42
1 93.30 96.55 76.57 83.99 91.29 70.91 73.09 89.40
2 89.76 96.31 57.79 65.99 94.42 66.34 48.34 95.88
3 70.16 50.38 56.23 76.06 83.06 60.09 54.54 80.98
4 44.32 80.63 47.60 44.77 68.39 50.12 49.37 73.66
5 91.49 96.51 74.48 89.11 86.46 77.84 66.47 81.63
6 60.98 81.63 61.97 76.29 75.58 61.89 69.84 87.02
7 76.80 83.98 71.37 77.69 86.56 77.49 56.51 89.57
8 89.77 77.10 85.53 84.42 93.33 92.96 78.87 98.11
9 84.40 98.48 82.64 93.03 97.14 81.02 65.62 97.20
10 82.47 91.91 61.00 86.16 87.58 46.08 39.27 86.42
11 90.77 99.32 81.41 91.18 94.78 88.12 70.29 93.19
12 67.64 91.80 52.87 63.56 85.45 60.61 55.59 87.75
13 85.99 50.80 45.48 35.34 29.04 12.91 32.89 91.60
14 93.52 54.56 15.81 22.79 94.12 38.21 15.10 98.49
15 36.62 30.93 37.71 43.70 42.78 40.65 29.70 28.96
16 95.62 65.11 90.13 94.97 97.47 95.32 98.27 99.01
17 60.54 95.74 56.24 91.71 78.72 59.11 13.29 96.44
18 63.92 17.39 58.47 70.21 54.61 47.93 10.01 79.70
19 62.12 79.06 49.36 65.77 81.07 51.19 26.84 84.13
Mean 75.04 75.88 60.45 71.29 79.94 61.26 49.17 86.33
Mean-Rank 4.45 3.20 6.20 4.25 3.00 5.70 7.05 2.15
Top-3 5 14 0 6 15 2 1 17
Top-1 0 8 0 1 1 0 0 10
Appendix F Extended experimental results
This section presents extended tables corresponding to the experiments described in Section 3.
F.1 How effective is ZEUS for clustering?
Tables 8, 9, and 10 contain the complete results for individual datasets from the experiments discussed in Section 3.2 in Tables 1 and 2, as presented in the main text. The reported values represent averages of ARI over 5 random seeds. In addition to the rows presenting outcomes for specific datasets, the tables include summary rows labeled Mean, Mean-Rank, Top-3, and Top-1. These represent, respectively: the average ARI, the average rank computed across all datasets, the number of times a given model appears in the top 3, and the total number of wins achieved by each method. It is worth noting that the Top-1 and Top-3 rows indicate clear wins and clear appearances in top-3 positions.
The overall conclusions align with those presented in the primary analysis. One noteworthy point is that ZEUS achieves strong performance in terms of the Top-3 and Top-1 statistics, being outperformed in this regard only in Table 9, where k-means and GMM score better. Another interesting observation is that GMM struggles with categorical features, which significantly worsens its average scores in Tables 9 and 10.
F.2 Are ZEUS’s assignments well calibrated?
Table 11, 12 and 13 contain detailed extension of Table 3
Table 11: Real/OpenML datasets - Brier
ID KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
14 0.7660 0.8579 0.8240 0.7144 0.5338 0.6409 0.7993 0.7006
15 0.0858 0.1548 0.2100 0.1791 0.1489 0.0644 0.2003 0.0973
16 0.5012 0.4982 0.4224 0.5372 0.3726 0.5020 0.6099 0.2761
18 0.6060 0.8529 0.7312 0.6961 0.6099 0.5354 0.6855 0.7094
22 0.8316 0.7533 1.0686 0.7684 0.5071 0.6496 0.7201 0.7173
35 0.0710 0.5016 0.4678 0.5711 0.5881 0.3987 0.3836 0.2552
51 0.3741 0.3741 0.3687 0.3844 0.3297 0.3345 0.3997 0.3333
53 0.4074 0.7052 0.3556 0.6285 0.3361 0.2979 0.5234 0.4000
56 0.2391 0.2344 0.2520 0.2703 0.2147 0.2088 0.2517 0.1829
61 0.2267 0.0570 0.3600 0.3796 0.3881 0.3445 0.4167 0.0979
187 0.0899 0.0622 0.1349 0.1146 0.3380 0.0785 0.2518 0.0762
377 0.8633 0.8320 0.6200 0.6275 0.5316 0.5047 0.5739 0.8508
458 0.0285 0.0214 0.5222 0.3407 0.3203 0.3656 0.4104 0.0048
481 0.8038 0.8134 0.4536 0.2136 0.3111 0.2146 0.4422 0.6986
694 1.0155 0.8831 0.9858 0.9342 0.7616 0.7525 0.7038 0.7621
721 0.3400 0.3400 0.6620 0.6031 0.4491 0.4359 0.7937 0.3400
733 0.1722 0.2859 0.3082 0.4937 0.2115 0.1129 0.3195 0.1244
745 0.2516 0.2516 0.1359 0.2766 0.3016 0.1793 0.2388 0.1384
756 0.2767 0.3615 0.2113 0.3262 0.3178 0.1329 0.4134 0.2767
796 0.0574 0.7943 0.1340 0.3088 0.2046 0.3002 0.5269 0.6253
820 0.2979 0.3720 0.3557 0.2876 0.2334 0.2594 0.3981 0.4596
840 0.8780 0.2790 0.2868 0.2625 0.2746 0.2445 0.4642 0.4665
854 0.6329 0.8203 0.3468 0.4481 0.4469 0.3372 0.3678 0.1266
1462 0.8484 0.8978 0.7688 0.8310 0.6674 0.7714 0.6747 0.0394
1495 0.0160 0.0080 0.1696 0.0397 0.2667 0.0713 0.1583 0.7200
1499 0.2190 0.2933 0.1867 0.1961 0.3939 0.1376 0.4063 0.1239
1510 0.1441 0.1161 0.2257 0.1954 0.2195 0.1432 0.2060 0.1371
1523 1.0516 0.7284 0.9019 0.9708 0.6099 0.7782 0.6316 1.0002
4153 0.7778 0.7178 0.6311 0.8111 0.5757 0.6377 0.8360 0.4651
40496 0.5064 0.9563 0.8312 0.8432 0.6733 0.6143 0.7376 1.0124
40682 0.2233 0.0747 0.2456 0.2251 0.2694 0.0685 0.3210 0.2884
40705 0.3879 0.3837 0.7095 0.4415 0.2541 0.3347 0.4304 0.3254
42261 0.2267 0.0570 0.3413 0.3760 0.3889 0.2854 0.3800 0.1081
42585 0.6279 0.9790 0.5558 0.6111 0.3494 0.7425 0.3795 0.0383
Mean 0.4366 0.4799 0.4643 0.4679 0.3941 0.3671 0.4722 0.3817
Mean-Rank 4.54 4.94 5.12 5.50 3.88 2.85 5.65 3.51
Top-3 12 9 8 4 19 22 6 21
Top-1 3 5 1 1 6 8 1 8
Table 12: Synthetic gaussian datasets - Brier
ID KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
0 0.1448 0.2200 0.1741 0.3558 0.4251 0.3264 0.3921 0.1242
1 0.1218 0.1886 0.2337 0.4718 0.4379 0.4109 0.4377 0.4502
2 0.1985 0.9985 0.2995 0.1922 0.4303 0.2321 0.5400 0.1952
3 0.0528 0.9078 0.0927 0.3012 0.6196 0.1628 0.4194 0.0413
4 0.1005 0.0575 0.1793 0.3351 0.4955 0.2312 0.3886 0.0792
5 0.1020 1.0950 0.1333 0.4892 0.3355 0.2059 0.3995 0.0881
6 0.0350 0.0156 0.1200 0.0489 0.2971 0.0770 0.3561 0.0209
7 0.0632 0.9781 0.1599 0.2610 0.4693 0.1207 0.6146 0.0543
8 0.1168 0.0844 0.2079 0.2786 0.4626 0.2970 0.5975 0.1631
9 0.0677 0.0092 0.3552 0.1925 0.2167 0.2090 0.4586 0.0573
10 0.1091 0.0292 0.1469 0.0460 0.3300 0.1229 0.3127 0.0655
11 0.0558 0.0117 0.0791 0.1370 0.4110 0.1122 0.4437 0.0234
12 0.1361 0.2296 0.4022 0.3492 0.5917 0.3230 0.3583 0.1545
13 0.0216 0.0115 0.0356 0.0629 0.3223 0.0980 0.1947 0.0270
14 0.1048 0.8000 0.3009 0.3776 0.3359 0.2380 0.3818 0.0626
15 0.0734 0.0278 0.2574 0.0746 0.3458 0.2483 0.3120 0.0352
16 0.0537 0.0448 0.0958 0.1747 0.1640 0.2528 0.1454 0.1770
17 0.1309 0.3513 0.4178 0.2891 0.4313 0.2756 0.5425 0.1308
18 0.1035 0.0452 0.1458 0.2319 0.5193 0.2117 0.2961 0.1939
19 0.1475 0.1132 0.3094 0.2989 0.2444 0.4612 0.3015 0.3942
Mean 0.0970 0.3110 0.2073 0.2484 0.3943 0.2308 0.3946 0.1269
Mean-Rank 2.30 3.25 4.55 4.95 6.65 4.85 6.75 2.70
Top-3 19 13 6 2 1 3 0 16
Top-1 2 11 0 1 0 0 0 6
Table 13: Synthetic transformed datasets - Brier
ID KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
0 0.5116 0.2135 0.6639 0.4457 0.5911 0.5063 0.7116 0.1282
1 0.0666 0.0484 0.2328 0.1514 0.4755 0.5045 0.2806 0.2718
2 0.0883 0.0330 0.4996 0.2822 0.4424 0.3441 0.5990 0.0394
3 0.3254 0.6286 0.5872 0.2118 0.3241 0.4862 0.7537 0.2678
4 0.6757 0.1547 0.5580 0.6123 0.3920 0.5158 0.3804 0.3190
5 0.0887 0.0359 0.3201 0.1517 0.6415 0.4114 0.3869 0.2802
6 0.5771 0.1788 0.4789 0.4049 0.5718 0.3069 0.5448 0.1763
7 0.3634 0.2875 0.4281 0.3326 0.5478 0.3379 0.6068 0.2461
8 0.1048 0.2738 0.1271 0.0817 0.4517 0.0784 0.3278 0.0167
9 0.3014 0.0120 0.3120 0.0940 0.4474 0.3329 0.2284 0.0481
10 0.1785 0.0763 0.3674 0.1249 0.3845 0.5057 0.5420 0.1387
11 0.1053 0.0130 0.1684 0.1128 0.6071 0.2580 0.3306 0.1768
12 0.3682 0.0657 0.6462 0.4740 0.3838 0.4585 0.6318 0.1936
13 0.0604 0.2549 0.4206 0.4292 0.3711 0.4073 0.3436 0.0355
14 0.0445 0.7982 1.1164 0.6259 0.3672 0.4949 0.6715 0.0097
15 0.7142 0.7088 0.6256 0.7924 0.5801 0.5741 0.4807 0.8854
16 0.0524 0.4745 0.1194 0.0303 0.5832 0.1743 0.1014 0.0809
17 0.2586 0.0230 0.2616 0.0394 0.3480 0.2156 0.6361 0.0177
18 0.2081 0.6698 0.3781 0.2409 0.3774 0.3861 0.7879 0.1056
19 0.5120 0.1821 0.7577 0.6415 0.3892 0.4845 0.5570 0.1542
Mean 0.2803 0.2566 0.4535 0.3140 0.4638 0.3892 0.4951 0.1796
Mean-Rank 3.95 3.00 5.85 3.95 5.75 5.00 6.15 2.35
Top-3 9 14 0 12 4 3 2 16
Top-1 0 8 0 2 0 0 1 9
F.3 How helpful is regularization for ZEUS?
Tables 14, 15 and 16 provide extended versions of Table 4 , which analyzes the impact of different combinations of regularization functions.
Table 14: Real/OpenML datasets - ablations ARI
ID $\mathcal{L}_{prob}$ $\mathcal{L}_{prob}+\mathcal{L}_{sep}$ $\mathcal{L}_{prob}+\mathcal{L}_{cp}$ $\mathcal{L}_{prob}+\mathcal{L}_{sep}+\mathcal{L}_{cp}$
14 48.40 43.49 46.44 50.56
15 86.64 87.17 82.94 81.28
16 69.87 73.33 81.87 74.03
18 43.06 50.80 49.62 51.63
22 63.27 56.86 65.58 56.05
35 52.33 74.58 76.90 85.12
51 32.31 31.54 41.84 43.66
53 5.17 27.38 25.85 35.76
56 58.48 61.34 20.84 66.41
61 48.93 88.57 65.37 85.15
187 88.38 91.50 88.22 88.19
377 46.31 45.16 54.56 55.30
458 98.16 98.49 98.49 99.19
481 2.77 25.87 1.53 8.68
694 42.76 48.86 30.35 33.38
721 43.28 5.39 43.28 43.28
733 34.35 24.94 76.83 74.97
745 59.56 22.80 4.46 73.85
756 76.15 63.49 38.42 50.31
796 6.56 3.96 14.52 13.99
820 39.95 39.95 34.72 28.87
840 0.95 4.13 31.14 27.66
854 12.37 12.37 24.32 76.14
1462 92.00 94.25 62.85 92.28
1495 -0.29 98.40 7.35 7.35
1499 60.63 44.86 71.34 82.40
1510 70.09 73.67 72.51 74.26
1523 25.68 28.13 17.38 19.60
4153 38.28 51.26 50.24 62.29
40496 34.15 33.63 30.68 32.25
40682 -0.39 23.38 42.79 53.15
40705 40.00 40.51 38.18 45.45
42261 48.93 88.57 65.37 85.15
42585 53.97 95.82 97.39 95.05
Mean 44.80 51.60 48.65 57.43
Mean-rank 3.00 2.37 2.68 1.96
Top-3 20 26 25 30
Top-1 2 9 6 15
Table 15: Synthetic gaussian datasets - ablations ARI
ID $\mathcal{L}_{prob}$ $\mathcal{L}_{prob}+\mathcal{L}_{sep}$ $\mathcal{L}_{prob}+\mathcal{L}_{cp}$ $\mathcal{L}_{prob}+\mathcal{L}_{sep}+\mathcal{L}_{cp}$
0 72.89 68.81 85.85 87.65
1 89.50 71.68 88.72 72.04
2 80.18 63.23 92.59 82.96
3 77.05 79.22 91.23 95.76
4 72.32 84.96 92.82 92.85
5 91.88 88.98 82.25 90.72
6 81.85 81.18 87.49 97.89
7 85.33 79.93 93.75 93.60
8 72.54 76.12 87.92 84.00
9 95.47 95.97 93.59 94.55
10 88.23 89.61 85.07 87.32
11 97.26 94.91 96.55 97.45
12 77.58 78.07 85.81 86.04
13 74.41 76.08 96.07 96.17
14 84.00 82.11 95.39 94.60
15 77.65 78.94 97.10 96.68
16 93.82 93.46 95.82 87.19
17 88.83 86.38 88.05 86.29
18 89.90 90.21 90.02 83.01
19 76.62 77.69 85.63 73.85
Mean 83.37 81.88 90.59 89.03
Mean-rank 2.80 3.00 2.00 2.20
Top-3 14 13 17 16
Top-1 3 3 7 7
Table 16: Synthetic transformed datasets - ablations ARI
ID $\mathcal{L}_{prob}$ $\mathcal{L}_{prob}+\mathcal{L}_{sep}$ $\mathcal{L}_{prob}+\mathcal{L}_{cp}$ $\mathcal{L}_{prob}+\mathcal{L}_{sep}+\mathcal{L}_{cp}$
0 67.29 71.62 85.76 87.42
1 91.98 91.46 96.88 89.40
2 93.17 93.84 92.36 95.88
3 68.26 71.68 80.53 80.98
4 80.93 81.64 84.22 73.66
5 81.38 80.06 95.99 81.63
6 84.29 82.06 80.67 87.02
7 83.35 73.07 96.83 89.57
8 97.42 98.61 97.16 98.11
9 97.63 96.43 96.37 97.20
10 91.32 77.37 90.58 86.42
11 91.84 89.07 99.17 93.19
12 74.93 92.20 91.14 87.75
13 0.32 0.73 90.67 91.60
14 98.00 98.83 98.16 98.49
15 43.28 48.54 30.50 28.96
16 93.34 90.38 98.11 99.01
17 97.62 98.21 95.86 96.44
18 82.64 82.22 86.94 79.70
19 77.93 67.71 83.78 84.13
Mean 79.85 79.29 88.58 86.33
Mean-rank 2.80 2.70 2.30 2.20
Top-3 15 14 15 16
Top-1 2 5 6 7
Appendix G Licensing and Third-Party Assets.
Our method builds upon the publicly available TabPFN codebase, which is released under the Apache 2.0 License. All real-world datasets used in the evaluation are sourced from OpenML.org, a platform hosting open datasets for machine learning research; all datasets used are publicly accessible and labeled as open data, although individual datasets may be subject to specific licenses (e.g., CC-BY or similar). We ensured that no proprietary or restricted datasets were used. Our released code is provided under the Apache 2.0 License and includes full instructions to reproduce the experiments.
Broader Impact.
Although ZEUS is designed as a general-purpose tool for clustering tabular data, its use in high-stakes domains like healthcare, finance, or criminal justice carries inherent risks. In such contexts, automated grouping of individuals—especially without labels or fairness constraints—can lead to biased or opaque outcomes. We therefore recommend that any deployment of ZEUS in sensitive applications be accompanied by fairness assessment and expert review.

### v2

@@
 However, these methods frequently suffer from instability due to their sensitivity to hyperparameter selection, a challenge exacerbated by the absence of supervised signals to guide optimization.
-Consequently, practitioners working with tabular data often resort to simpler, classical algorithms like k-means, despite their limited capacity for capturing complex underlying data structures, simply to avoid extensive manual tuning.
+Consequently, practitioners working with tabular data often resort to simpler, classical algorithms like k-means macqueen1967some , despite their limited capacity for capturing complex underlying data structures, simply to avoid extensive manual tuning.
 We address these challenges with ZEUS – a zero-shot transformer-based model for embedding new tabular datasets in a form convenient for unsupervised separation (=clustering), without the need for additional fine-tuning.
 Given a new dataset, ZEUS returns its transformed representation, where clusters can be easily discovered using simple methods, like k-means.
-Since it works as a zero-shot learner, it significantly reduces hyperparameter tuning complexity and the computation time, enabling effective clustering in seconds.
+Since it works as a zero-shot learner, it significantly reduces hyperparameter tuning complexity and computation time, enabling effective clustering in seconds.
 It is a plug-and-play solution that requires no fine-tuning and runs in a single forward pass for new datasets.
 Inspiration for ZEUS stems from the Prior-data Fitted Networks (PFNs) tabpfn , a recently introduced framework highlighting the potential of in-context learning for tabular data.
-PFNs operate by feeding a transformer contexts representing complete tasks, i.e., training data along with query samples for which the model predicts labels.
+PFNs operate by feeding a transformer transformer contexts representing complete tasks, i.e., training data along with query samples for which the model predicts labels.
 Despite their impressive performance, they are limited to supervised problems.
@@
 Additionally, in contrast to TabPFN, ZEUS is a zero-shot model and during inference does not rely on any context labels.
-Unlike traditional methods that optimize arbitrary heuristics (e.g., DEC), ZEUS approaches clustering by learning how to invert data generation processes.
+Unlike traditional methods that optimize arbitrary heuristics (e.g., DEC xie2016unsuperviseddeepembeddingclustering ), ZEUS approaches clustering by learning how to invert data generation processes.
 In particular, it is pre-trained on synthetic datasets to learn how to infer cluster assignments.
 The datasets have known latent structures, which enables supervised guidance for the training, and by generating diverse datasets we supply it with the prior knowledge about what can constitute a possible clustering structure, enabling effective generalization to new real-world datasets during inference.
-Figure 1 illustrates the key concepts of the method while Figure 2 presents representations generated by ZEUS.
+Figure 1 illustrates the key concepts of the method while Figure˜2 presents representations generated by ZEUS.
 Experiments demonstrate that ZEUS consistently matches or surpasses both classical clustering algorithms and recent deep learning-based approaches, establishing it as a powerful and practical tool for the unsupervised analysis of tabular data.
-Beyond its strong empirical performance, we also provide theoretical justification by showing that ZEUS fits the framework of Prior-data Fitted Networks, thereby reinforcing its theoretical soundness (see Section 2.3).
-We supplement the paper with an appendix that includes background on prior-data fitted networks, details of the synthetic data generation process, and additional details of the experimental results.
-The code used in this paper is available at https://github.com/gmum/zeus
+Beyond its strong empirical performance, we also provide theoretical justification by showing that ZEUS fits the framework of Prior-data Fitted Networks, thereby reinforcing its theoretical soundness (see Section˜2.3).
+We supplement the paper with an appendix that includes background on prior-data fitted networks, specifics of the synthetic data generation process, details of the experimental results, and additional experimental studies.
+The code used in this paper is available at https://github.com/gmum/zeus.
 Figure 1: Schematic characterization of ZEUS: (left) synthetic datasets generation; (middle) pre-training on datasets with known labels; (right) deployment of a frozen model for real-world tasks.
@@
 Traditional machine learning models $f_{\theta}$ are trained by optimizing parameters $\theta$ to solve a specific fixed task, such as classification or clustering, over a fixed dataset.
-On the other hand, in-context learning (ICL) 10.5555/3600270.3602487 ; lee2023supervised ; han-etal-2023-understanding enables pre-trained models $f_{\theta^{\ast}}$ to adapt to new tasks or datasets without ever updating the pre-trained parameters $\theta^{*}$.
+On the other hand, in-context learning (ICL) garg2023transformerslearnincontextcase ; lee2023supervised ; han-etal-2023-understanding enables pre-trained models $f_{\theta^{\ast}}$ to adapt to new tasks or datasets without ever updating the pre-trained parameters $\theta^{*}$.
 Similarly, zero-shot learning allows models to address problems they were never explicitly trained on, without requiring any examples of the target task.
@@
 The pre-training of zero-shot models typically involves massive and diverse datasets that encourage the model to learn rich, generalizable representations of underlying patterns and relationships.
-This usually is done using transformers, as they both can process set-valued inputs (e.g., accept a whole dataset at once) and model complex dependencies.
+This is usually done using transformers, as they both can process set-valued inputs (e.g., accept a whole dataset at once) and model complex dependencies.
 Our approach involves pre-training a transformer $f_{\theta}$ on a diverse collection of datasets $\mathcal{D}$ such that $\mathcal{D}:=(\mathbf{x},\mathbf{y})$, where $\mathbf{y}=\{y_{i}\}$ denotes ground-truth labels (e.g., clusters) and $\mathbf{x}=\{x_{i}\}$ are feature vectors.
@@
 $\mathcal{L}_{prob}=-\sum_{i}\log p_{y_{i}}(x_{i}).$ (2)
-For specifying $p_{y}(x)$ we draw an inspiration from Gaussian Mixture Models (GMMs) dempster1977maximum ; hansen1982large , where each cluster $k$ is represented by a Gaussian distribution centered at ${c}_{k}$, and data points are probabilistically assigned to clusters through probabilities $p_{k}(x_{i})$, based on their proximity to these centroids.
+For specifying $p_{y}(x)$, we draw inspiration from Gaussian Mixture Models (GMMs) dempster1977maximum ; hansen1982large , where each cluster $k$ is represented by a Gaussian distribution centered at ${c}_{k}$, and data points are probabilistically assigned to clusters through probabilities $p_{k}(x_{i})$, based on their proximity to these centroids.
 Consequently, alongside the encoded representations, we need to define a set of $K$ cluster centroids (e.g., prototypical representations) $\{{c}_{k}\}_{k=1}^{K}$, which reside in the same representation space $S$.
 We then jointly optimize $\mathcal{L}$ w.r.t $\theta$ and $\{{c}_{k}\}$ such that the representations ${z}(x_{i})$ become structured in a way undisclosing underlying probabilistic structures in datasets.
-We aim at higher probabilities $p$ for points closer these centroids.
-To achieve this, we first define scores $\alpha_{k}(x_{i})=-\|{z}(x_{i})-{c}_{k}\|^{2}$ to quantify the compatibility between representations and centroids ${c}_{k}$, and then transform them into cluster membership probabilities using softmax, yielding soft cluster assignments $p_{k}(x_{i})=\frac{\exp(\alpha_{k}(x_{i}))}{\sum_{j=1}^{K}\exp(\alpha_{j}(x_{i}))}.$ Note that the ground-truth cluster assignments $y_{i}$ are known during pre-training, and hence, the maximum likelihood centroids can be straightforward estimated as $\hat{c}_{k}=\frac{1}{N_{k}}\sum_{\{i:y_{i}=k\}}{z}(x_{i}),$ where $N_{k}$ denotes the number of data points assigned to the $k$-th cluster.
-Following this formulation, the optimization needs to performed only w.r.t the parameters $\theta$ of the neural network $f_{\theta}$, as the centroids are already specified by them.
+We aim for higher probabilities $p$ for points closer to these centroids.
+To achieve this, we first define scores $\alpha_{k}(x_{i})=-\|{z}(x_{i})-{c}_{k}\|^{2}$ to quantify the compatibility between representations and centroids ${c}_{k}$, and then transform them into cluster membership probabilities using softmax, yielding soft cluster assignments $p_{k}(x_{i})=\frac{\exp(\alpha_{k}(x_{i}))}{\sum_{j=1}^{K}\exp(\alpha_{j}(x_{i}))}.$ Note that the ground-truth cluster assignments $y_{i}$ are known during pre-training, and hence, the maximum likelihood centroids can be straightforwardly estimated as $\hat{c}_{k}=\frac{1}{N_{k}}\sum_{\{i:y_{i}=k\}}{z}(x_{i}),$ where $N_{k}$ denotes the number of data points assigned to the $k$-th cluster.
+Following this formulation, the optimization needs to be performed only w.r.t the parameters $\theta$ of the neural network $f_{\theta}$, as the centroids are already specified by them.
 Besides the centroids, we additionally consider another concept from GMMs, namely cluster priors ${\pi}_{k}$.
@@
 Inference.
-The structural relations learned during pre-training are leveraged during inference when the representations ${z}$ produced by the model $f_{\theta^{\ast}}$ are used to create a predictive distribution for new, previously unseen inputs $\mathbf{x}^{*}$ as $\hat{\mathbf{p}}^{*}:=p\bigl{(}\mathbf{y}^{*}\mid f_{\theta^{\ast}}(\mathbf{x}^{*})\bigr{)}.$ At this point, the labels $\mathbf{y}^{*}$ are unknown and therefore, cannot be used to estimate the centroiods nor the priors.
+The structural relations learned during pre-training are leveraged during inference when the representations z{z} produced by the model $f_{\theta^{\ast}}$ are used to create a predictive distribution for new, previously unseen inputs $\mathbf{x}^{*}$ as $\hat{\mathbf{p}}^{*}:=p\bigl(\mathbf{y}^{*}\mid f_{\theta^{\ast}}(\mathbf{x}^{*})\bigr).$ At this point, the labels $\mathbf{y}^{*}$ are unknown and therefore cannot be used to estimate the centroids nor the priors.
 Hence, to structure the obtained representation into clusters, we use a traditional learning algorithm, for example, similar to GMMs relying on Expectation-Maximization or simply k-means.
 Regularization.
-Although theoretically $\mathcal{L}_{prob}$ is sufficient to build a clustering representation, we experimentally verified that the introducing additional regularizers further improves the predictions in the inference (see Section 3.4).
+Although theoretically $\mathcal{L}_{prob}$ is sufficient to build a clustering representation, we experimentally verified that the introduction of additional regularizers further improves the predictions in the inference (see Section˜3.4).
 First, we aim to ensure that the representations ${z}(x_{i})$ associated with a particular cluster are compactly distributed around their corresponding centroid to enhance intra-cluster cohesion.
-A point concentration regularizer that explicitly minimizes the distance between representations and the centroids, analogous to the k-means objective of minimizing within-cluster sum of squares, we define as
+We define a point concentration regularizer that explicitly minimizes the distance between representations and the centroids, analogous to the k-means objective of minimizing within-cluster sum of squares:
 $\mathcal{L}_{cp}=\sum_{k}\sum_{i:y_{i}=k}\alpha_{k}(x_{i}).$ (4)
@@
 In particular, we construct it as a latent variable model (LVM), assuming that each dataset $\mathcal{D}$ is sampled from a $K$-component mixture of distributions.
-Formally, the probability of a data point $x\in\mathcal{D}$ under this model is $p(x)=\sum_{k=1}^{K}p(y=k)\,p(x\mid y=k),$ where $p(y)$ is a categorical distribution, and $p(x\mid y=k)$ represents the $k$-th (continuous) component.
+Formally, the probability of a data point $x\in\mathcal{D}$ under this model is
+$p(x)=\sum_{k=1}^{K}p(y=k)\,p(x\mid y=k),$
+where $p(y)$ is a categorical distribution, and $p(x\mid y=k)$ represents the $k$-th (continuous) component.
 Although the categories $y$ remain latent for real datasets, they are known during synthetic data generation and can serve as ground-truth labels – a property integral to pre-training ZEUS.
+This procedure, coupled with Theorem 1 (in Appendix) ensures that a sufficiently expressive model trained on synthetic datasets can handle real datasets drawn from arbitrary mixtures of distributions.
 The number of categories $K$ we sample uniformly between 2 and 10, and the observations $x$ are sampled from multivariate Gaussian distributions.
-To control the complexity of datasets, we introduce a constraint that provides a sufficient separation between each pair of components.
+To control the complexity of datasets, we introduce a constraint (Eq.
+7 in Appendix) that ensures a sufficient separation between each pair of components.
 For each component, we generate between $50$ and $800$ samples.
@@
 Finally, we append a certain fraction of categorical features in one-hot encoding scheme to selected datasets, which is followed by an optional PCA reduction to keep the data dimension at the requested maximum level.
-Complete details of the prior-data generating process can be found in Appendix B.
+Complete details of the prior-data generating process can be found in Appendix B.2.
+Theoretical justification of this procedure is presented in Appendix B.1.
 The above strategy for generating clustering datasets illustrates the key idea behind ZEUS.
@@
 By framing the approach described in Section 2.1 as a Prior-Data Fitted Network (PFN), we show ZEUS implicitly performs approximate Bayesian averaging.
-[…]
-ZEUS instead of directly outputting $q_{\theta}(y|x,\mathcal{D}_{ctx})$, maps inputs $x$ to latent representation vectors ${z}(x)$.
+Given a prior distribution $p(\mathcal{D})$ over (synthetic) datasets $\mathcal{D}$, a PFN parameterized by $\theta$ is trained to minimize the negative log-likelihood of predicting held-out labels within datasets sampled from this prior.
+The associated loss function is defined as: $\mathcal{L}_{\textit{PFN}}(\theta)=\mathbb{E}_{\mathcal{D}_{ctx}\cup\{(x,y)\}\sim p(\mathcal{D})}[-\log q_{\theta}(y|x,\mathcal{D}_{ctx})].$ Minimizing the Prior-Data Negative Log-Likelihood is then equivalent to minimizing the expected Kullback–Leibler divergence between the network’s predictive distribution and the true Posterior Predictive Distribution (PPD) $p(y|x,\mathcal{D}_{ctx})=\int_{\Phi}p(y|x,\phi)p(\mathcal{D}_{ctx}|\phi)p(\phi)d\phi$ (see Corollary 1.1 in bayes_in_context ).
+A pre-trained PFN approximates this integral implicitly, yielding a distribution $q_{\theta}(y|x,\mathcal{D}_{ctx})$ directly from forward propagation of the network.
+In Appendix˜A, a more detailed explanation of PFNs was provided.
+ZEUS instead of directly outputting $q_{\theta}(y|x,\mathcal{D}_{ctx})$, maps inputs $x$ to latent representation vectors z(x){z}(x).
 The probabilistic assignments $p_{y}(x)$ are then constructed from these vectors according to Eq.
@@
 This structure might be suboptimal for classification tasks and the true PPD may not belong to the family of attainable solutions.
-As explained above our clustering-based extension is theoretically sound, nonetheless, the assumption could limit representational expressivity compared to the more flexible transformers employed for the original PFNs tabpfn .
+As explained above, our clustering-based extension is theoretically sound, nonetheless, the assumption could limit representational expressivity compared to the more flexible transformers employed for the original PFNs tabpfn .
 On the other hand, the enforced structures shall be more appropriate for unsupervised tasks.
 3 Experiments
-This section provides an experimental evaluation of the clustering performance of our method.
+This section presents an experimental evaluation of the clustering performance of our method. Due to space constraints, further results are provided in Appendix G.
 3.1 Experimental setup
 Model architecture: ZEUS relies on a transformer architecture similar to TabPFN (tabpfn, ).
-It consists of 12 attention blocks, each with 6 heads and a token dimension of 512, employing GeLU as the activation function.
-Following the TabPFN approach, each data point is linearly embedded prior to the transformer component and then treated as a token.
+It consists of 12 attention blocks, each with 6 heads and a token dimension of 512, with GeLU activation employed.
+Following the TabPFN design, each data point is first embedded using a linear transformation and then passed as a token to the transformer.
+For the unsupervised setting, no label embeddings are ever created or presented to the model.
+Similarly, for the zero-shot case, no query set is used either, and consequently, attention is computed solely over the support set.
+Finally, unlike TabPFN, we omit any additional MLP decoder after the transformer.
 Pre-training: In the pre-training phase, we sample datasets from the mixture of Gaussians and transformed mixtures in equal proportions (1:1), generating 1000 unique dataset batch samples for each epoch.
 For training, we employ the Adam optimizer along with a cosine learning rate scheduler with warm-up, using a learning rate of 2e-5.
-The plot illustrating model improvements during the pre-training process is available in Appendix C.
-Inference: During inference, preprocessing of each dataset involves standardizing numerical features, followed by scaling them to the range $[-1,1]$, whereas categorical features are transformed using one-hot encoding.
+The plot illustrating model improvements during the pre-training process is available in Appendix˜C.
+Inference: During inference, preprocessing of each dataset involves standardizing numerical features, followed by scaling them to the range [1,1][-1,1], whereas categorical features are transformed using one-hot encoding.
 The input size of our model is fixed to 30.
 For datasets with lower dimensionality, we pad the missing positions with zeros, while for higher-dimensional datasets, we reduce the number of input features via Principal Component Analysis (PCA).
-Unless stated otherwise, at inference we use k-means applied to the normalized (=scaled to $[-1,1]$) transformer output in order to obtain clusters from our learned representation.
-Datasets: For evaluation, we consider three groups of datasets: real datasets from OpenML (Real), synthetic mixtures of Gaussians (Syn.
+Unless stated otherwise, at inference we use k-means applied to the normalized (=scaled to [1,1][-1,1]) transformer output in order to obtain clusters from our learned representation.
+Datasets: For evaluation, we consider three groups of datasets: real datasets from OpenML OpenML2025 (Real), synthetic mixtures of Gaussians (Syn.
 Gauss.), and synthetic mixtures of Gaussians transformed by ResNet-like neural networks (Syn.
@@
 Both types of synthetic datasets are augmented with categorical variables.
-The process of generating synthetic datasets is described in Section 2.2.
+The process of generating synthetic datasets is described in Section˜2.2.
 Each dataset contains at most 2000 samples as per model design and due to memory limitations.
@@
 Baselines: We compare ZEUS against a wide spectrum of state-of-the-art clustering methods used for tabular data.
-It includes $k$-means (KM), Gaussian Mixture Model (GMM), and deep-learning methods based on autoencoder (AE) architectures, including DEC xie2016unsuperviseddeepembeddingclustering , IDEC guo_improved_2017 , and G-CEALS rabbani2025deep .
-We additionally consider also $k$-means and GMM applied in the autoencoder latent space (respectively referred to as AE-KM and AE-GMM) to compare the clustering quality of our representation against the one obtained from the autoencoder.
+It includes $k$-means (KM), Gaussian Mixture Model (GMM), and deep-learning methods based on autoencoder (AE) architectures, including DEC xie2016unsuperviseddeepembeddingclustering , IDEC guo_improved_2017 , IDC idc , and G-CEALS rabbani2025deep .
+We additionally consider $k$-means and GMM performed in the autoencoder latent space (respectively referred to as AE-KM and AE-GMM), along with $k$-means applied to representations obtained from TabPFN-Unsupervised hollmann2025accurate (TabPFN) and SCARF scarf , to assess the clustering quality of different feature representations.
 For AE-based baselines, we employ the standard configuration used in the prior literature, i.e., an architecture comprising of hidden layers with sizes [500, 500, 2000].
@@
 Gauss, Syn.
-Transf.), while detailed results for individual datasets can be found in Appendix F.
+Transf.), while detailed results for individual datasets can be found in Appendix˜F.
 For the reader’s convenience, we bold the best results and underline the second-best ones.
 3.2 How effective is ZEUS for clustering?
-To evaluate performance of the clustering methods, we employ a standard evaluation procedure, where clusters identified by models are expected to correspond to undisclosed ground-truth classes.
+To evaluate the performance of the clustering methods, we employ a standard evaluation procedure, where clusters identified by models are expected to correspond to undisclosed ground-truth classes.
 We use the Adjusted Rand Index (ARI) for quantitative evaluation, and to improve readability, we scale the ARI values by a factor of 100, where 100 indicates perfect clustering and values near 0 represent random grouping.
@@
 Table 1: Clustering quality (ARI) of ZEUS versus competing methods (higher is better).
-KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
-Real 55.54 48.49 51.43 53.56 55.93 54.57 40.37 57.43
-Syn. Gauss. 89.90 76.93 81.26 81.40 89.36 82.57 62.84 89.03
-Syn. Transf. 75.04 75.88 60.45 71.29 79.94 61.26 49.17 86.33
+KM GMM AE-KM AE-GMM DEC IDEC IDC G-CEALS TabPFN SCARF ZEUS
+Real 55.54 48.49 51.43 53.56 55.93 54.57 52.28 40.37 31.32 26.95 57.43
+Syn. Gauss. 89.90 76.93 81.26 81.40 89.35 82.57 66.43 62.84 55.97 8.32 89.03
+Syn. Transf. 75.04 75.88 60.45 71.29 79.94 61.26 66.78 49.17 15.66 2.48 86.33
 Table 2: Average rank of the methods used in the benchmark (lower is better).
-KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
-Real 3.90 4.57 4.72 4.29 3.96 4.31 6.65 3.60
-Syn. Gauss. 2.60 3.35 5.40 5.45 3.10 5.50 7.70 2.90
-Syn. Transf. 4.45 3.20 6.20 4.25 3.00 5.70 7.05 2.15
+KM GMM AE-KM AE-GMM DEC IDEC IDC G-CEALS TabPFN SCARF ZEUS
+Real 4.69 5.65 5.72 5.24 4.69 5.01 5.62 8.18 8.22 8.85 4.13
+Syn. Gauss. 2.65 3.65 5.65 5.60 3.23 5.70 7.95 8.90 8.75 11.00 2.92
+Syn. Transf. 4.80 3.50 6.85 4.50 3.20 6.35 6.05 7.80 9.75 11.00 2.20
+Figure 3: Clustering time vs. input size.
 We observe that ZEUS achieves competitive performance for all three groups of datasets.
@@
 In terms of average rank it gets the second place.
-Figure 3: Clustering time vs. input size.
 Among the AE-based methods, DEC performs the best, while G-CEALS consistently demonstrates the weakest performance across all datasets.
 K-means and GMM, as representatives of the classical methods, perform reasonably well.
-In particular, k-means consistently lands among the best performing methods, which justifies its broad adoption among practitioners, despite being one of the most basic approaches.
-The detailed results including scores for individuals datasets corresponding to the averages in Tables 1 and 2 can be found in Appendix F.1
-Finally, Figure 3 illustrates how the examined clustering methods scale with the increasing number of input data points.
+In particular, k-means consistently lands among the best-performing methods, which justifies its broad adoption among practitioners, despite being one of the most basic approaches.
+On the other hand, approaches such as SCARF or TabPFN are well-suited for tasks like classification, thanks to their strong performance in feature extraction.
+However, without additional regularization during training or specialized post-hoc fine-tuning (such as with DEC), effective separation of their representations is difficult.
+The detailed results, including scores for individual datasets corresponding to the averages in Tables 1 and 2, can be found in Appendix F.1
+Finally, Figure˜3 illustrates how the examined clustering methods scale with the increasing number of input data points.
 ZEUS maintains almost constant time while being only slightly slower than the basic k-means.
@@
 Table 3 reports the average Brier score computed over 5 seeds for all datasets in respective groups.
-For the analysis, we used the vanilla variant of ZEUS without regularization, which is paired with the GMM clustering since k-means does not provide soft assignments.
+For the analysis, we used the vanilla variant of ZEUS, which is paired with the GMM clustering since k-means does not provide soft assignments.
 The covariances were constrained to be identities as implied by Eq.
@@
 ZEUS demonstrates outstanding performance, achieving the best results for the Synthetic Transformed datasets while ranking second on both the OpenML and Synthetic Gaussian collections.
-The benefits of using ZEUS representations are especially visible in comparison against the vanilla GMM.
+The benefits of using ZEUS representations are especially visible in comparison with the vanilla GMM.
 Although the GMM achieves the second-best score on transformed data, its performance across the remaining datasets is merely modest.
@@
 Table 4 shows that a model pre-trained using only $\mathcal{L}_{prob}$ and its variant with $\mathcal{L}_{sep}$ included are insufficient for properly separating the transformer’s representations for the clustering task, as they both struggle with clustering the synthetic datasets encountered during pre-training.
-On the other hand, the results for $\mathcal{L}_{prob}+\mathcal{L}_{cp}$ imply a positive impact by the compact loss component, $\mathcal{L}_{cp}$, especially for synthetic data, which however does not translate to strong performance on the OpenML benchmark.
+On the other hand, the results for $\mathcal{L}_{prob}+\mathcal{L}_{cp}$ imply a positive impact by the compact loss component, $\mathcal{L}_{cp}$, especially for synthetic data, which, however, does not translate to strong performance on the OpenML benchmark.
 We conclude that all three loss components are necessary for good performance, with $\mathcal{L}_{cp}$ being crucial for pre-training and $\mathcal{L}_{sep}$ important for real data.
@@
 Hence, for future work, one may want to explore alternative data-generating priors that may be more appropriate for these outliers.
+3.5 What data-generating prior is optimal?
+As mentioned in Section 2.2, constructing an appropriate data-generating prior $p(D)$ is crucial for maximizing performance and generalizability of ZEUS.
+To validate this claim, we conduct an ablation study evaluating four prior designs: Gaussian + Categorical, NN-transformed + Categorical, Gaussian + NN-transformed, and Gaussian + NN-transformed + Categorical (the standard ZEUS model).
+Each configuration name directly reflects the combination of priors used for pre-training (see Section B.2 in the Appendix for further details).
+For a fair comparison, we keep the default hyperparameter setup unchanged, except for the prior itself.
+Table 5: ARI scores (averages; in rows) for ZEUS variants pre-trained on different priors (columns).
+Gauss. + Cat. NN-transf. + Cat. Gauss. + NN-transf. Gauss. + NN-transf. + Cat.
+Real 40.59 50.90 52.00 57.43
+Syn. Gauss. 92.61 89.90 75.25 89.03
+Syn. Transf. 73.34 87.04 71.29 86.33
+The results of the study are summarized in Table 5.
+As expected, the Gaussian + Categorical model achieves top performance on the Synthetic Gaussian datasets, while the NN-transformed + Categorical model excels for the Synthetic Transformed data.
+Notably, the latter also performs well on the Gaussian sets, as its prior is built upon transformed Gaussian mixtures.
+In contrast, the Gaussian + NN-transformed model yields relatively poor average ARI scores across both synthetic benchmarks.
+This drop is primarily caused by a significant decline in ARI performance on the categorical datasets.
+Nonetheless, its average rank remains consistently competitive.
+Finally, for the OpenML datasets, the best performance is achieved when pre-training includes all three types of priors, underscoring the complementary importance of each prior in the training process.
 4 Related work
@@
 A pioneering Deep Embedded Clustering (DEC) xie2016unsuperviseddeepembeddingclustering improves the target data representation by training the autoencoder and computing soft assignments in its latent space via Student’s $t$-distribution.
-In a concurrent work yang2017towards , the authors perform joint dimensionality reduction using AE and k-means clustering in the latent space of AE.
-Improved DEC (IDEC) guo_improved_2017 extends DEC by jointly optimizing reconstruction and clustering objectives, while spectral variations replace k-means steps with graph-based updates 8970928 .
-G-CEALS rabbani2025deep replaces the $t$–distribution assumption with multivariate Gaussian clusters.
-To eliminate the assumption of explicit distribution prior, DEPICT dizaji2017deepclusteringjointconvolutional attaches a softmax layer on top of an embedding network and train with cross-entropy loss.
-Finally, IDC svirsky2023interpretable predicts interpretable cluster assignments at the instance and cluster levels.
-However, most of these deep learning models require careful hyperparameter tuning and early stopping which is unrealistic in the fully unsupervised setting due to the lack of labels 9506051 ; sadeghi2024deepclusteringselfsupervisionusing .
-Moreover, the optimization process has to be performed on each dataset, which is often time-consuming.
-Although multiple deep clustering approaches are currently in use, most of them are designed for texts or images znalezniak2023contrastive ; cai2023semantic ; li2024image and cannot be directly adapted to tabular data due to the lack of a dominant neural architecture for heterogeneous tabular inputs grinsztajn2022tree ; mcelfresh2023neural .
+In a concurrent work yang2017towards , the authors perform joint dimensionality reduction using AE and k-means clustering in the latent space.
+Then, improved DEC (IDEC) guo_improved_2017 extends DEC by jointly optimizing reconstruction and clustering objectives, while spectral variants replace k-means steps with graph-based updates 8970928 .
+G-CEALS rabbani2025deep replaces the tt–distribution assumption with multivariate Gaussian clusters.
+DEPICT dizaji2017deepclusteringjointconvolutional attaches a softmax layer on top of an embedding network and trains with cross-entropy loss to eliminate the assumption of an explicit distribution prior.
+Finally, IDC idc predicts interpretable cluster assignments at the instance and cluster levels.
+Most of these deep learning models require careful hyperparameter tuning and early stopping which is unrealistic in the fully unsupervised setting due to the lack of labels 9506051 ; sadeghi2024deepclusteringselfsupervisionusing .
+Moreover, the optimization process has to be performed for each dataset, which is often time-consuming.
+Although multiple deep clustering approaches are currently in use, most of them are designed for texts or images znalezniak2023contrastive ; cai2023semantic ; li2024image and cannot be directly adapted for tabular data due to the lack of a dominant neural architecture for heterogeneous tabular inputs grinsztajn2022tree ; mcelfresh2023neural .
 Representation learning for tabular data.
 Self-supervised learning (SSL) has been transformative in domains like vision and language chen2020simple ; grill2020bootstrap but has struggled to show similar success in tabular data yoon2020vime ; zhu2023xtab .
-Large diversity of data and lack of pre-defined correlation between features prevents from designing universal pretext tasks or augmentations as well as transferability between domains zhao2024comparison .
-In-context learning seems to be a promising direction in representation learning, which enables a single pre-trained model to adapt to new tasks in a zero-shot setting brown2020languagemodelsfewshotlearners .
-TabPFN treats small tabular datasets as contexts of (features, labels) and achieves state-of-the-art classification in one forward pass tabpfn .
+A large diversity of data and a lack of pre-defined correlation between features hinder the design of universal pretext tasks or augmentations, as well as the transfer learning between domains zhao2024comparison .
+On the other hand, in-context and zero-shot learning brown2020languagemodelsfewshotlearners enable the use of a single pre-trained model for multiple tasks out of the box, without any additional tuning.
+In particular, TabPFN treats small tabular datasets as contexts consisting of features along with labels, and achieves state-of-the-art classification in a single forward pass tabpfn .
 Although theoretical analyses reveal that transformers can implicitly implement algorithms such as gradient descent in context, their use is currently restricted to supervised problems garg2023transformerslearnincontextcase .
@@
 Acknowledgments and Disclosure of Funding
+We are grateful to the Reviewers for their effort and insightful comments on the paper.
 This research is part of the project No.
@@
 [Uncaptioned image] [Uncaptioned image]
-The research of M. Śmieja and P. Marszałek was supported by the National Science Centre (Poland), grant no. 2023/50/E/ST6/00169.
-The research of J. Tabor was supported by the National Science Centre (Poland), grant no. 2023/49/B/ST6/01137.
-Witold Wydmański is supported by the Ministry of Science grant no. PN/01/0195/2022 and NCN Sonata BIS grant number 2020/38/E/NZ2/00598.
+The research of P.
+Marszałek and M.
+Śmieja was supported by the National Science Centre (Poland), grant no.
+2023/50/E/ST6/00169.
+The research of Witold Wydmański was supported by the Ministry of Science grant no.
+PN/01/0195/2022 and NCN Sonata BIS grant number 2020/38/E/NZ2/00598.
+The research of J.
+Tabor was supported by the National Science Centre (Poland), grant no.
+2023/49/B/ST6/01137.
+Some experiments were performed on servers purchased with funds from the flagship project entitled “Artificial Intelligence Computing Center Core Facility” from the DigiWorld Priority Research Area within the Excellence Initiative – Research University program at Jagiellonian University in Krakow.
+We also gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no.
+PLG/2024/017893.
 Appendix A Background on Prior-Data Fitted Networks
@@
 Given a training dataset $\mathcal{D}_{ctx}=\{(x_{i},y_{i})\}_{i=1}^{n}$, a new input $x$, and a prior over hypotheses $\phi$, the Bayesian posterior predictive distribution is formally expressed as:
-[…]
+$p(y|x,\mathcal{D}_{ctx})=\int_{\Phi}p(y|x,\phi)p(\mathcal{D}_{ctx}|\phi)p(\phi)d\phi,$
 integrating over all hypotheses $\phi\in\Phi$.
@@
 Appendix B Details of synthetic data generation process
-We use three types of probabilistic models to generate data from a category:
+B.1 Theory for synthetic-to-real data generalization
+Theorem 1 (Univeral Approximation Theorem for Mixture Distributions).
+Let $P=\sum_{i=1}^{k}\pi_{i}P_{i}$ be a mixture of distributions on $\mathbb{R}^{d}$, where each $P_{i}$ is a probability distribution, and $\pi_{i}>0$, $\sum_{i=1}^{k}\pi_{i}=1$.
+Then there exists a mixture of Gaussians $Q=\sum_{i=1}^{k}\pi_{i}\mathcal{N}(\mu_{i},\Sigma_{i})$ and a neural network $F$ such that we can approximate $P$ with arbitrarily small error by the pushforward measure $F_{\#}Q$.
+Additionally, we can select $F$ so that $F_{\#}N(\mu_{i},\Sigma_{i})$ approximates $P_{i}$ with arbitrarily small error.
+Proof.
+We proceed in several steps.
+Step 1: Approximation of individual components.
+Making use of the universal approximation theorem, for each component $P_{i}$ there exists a neural network $F_{i}$ which transforms an arbitrary Gaussian $Q_{i}=\mathcal{N}(\mu_{i},\Sigma_{i})$ into $P_{i}$, i.e.
+$F_{i\#}Q_{i}\approx P_{i}$.
+Step 2: Constructing a unified neural network.
+We now define a single neural network $F$ that can represent all the $F_{i}$ networks.
+Define $F:\mathbb{R}^{d}\times\{1,\dots,k\}\to\mathbb{R}^{d}$ such that $F(z,i)\approx F_{i}(z)$.
+This can be implemented by encoding the index $i$ as a one-hot vector or embedding and concatenating it to $z$, enabling $F$ to learn the behavior of each $F_{i}$ conditioned on the index $i$.
+The network $F$ can be extended to the domain $\mathbb{R}^{d}\times\mathbb{R}$ in an arbitrary way.
+Step 3: Approximation of the full mixture.
+Since each $F_{\#}\mathcal{N}(\mu_{i},\Sigma_{i})$ approximates $P_{i}$, the mixture $F_{\#}Q=\sum_{i=1}^{k}\pi_{i}F_{\#}\mathcal{N}(\mu_{i},\Sigma_{i})$ approximates $P=\sum_{i=1}^{k}\pi_{i}P_{i}$.
+Thus, we have constructed a mixture of Gaussians $Q$ and a neural network $F$ such that $F_{\#}Q$ approximates $P$, and each component $F_{\#}\mathcal{N}(\mu_{i},\Sigma_{i})$ approximates $P_{i}$ arbitrarily well.
+This completes the proof.
+∎
+B.2 Insights into data-generating priors
+We use three types of probabilistic procedures to generate the prior data:
   1. 1.
     Gaussian.
-The continuous features of each cluster follow a multivariate Gaussian distribution with carefully constructed means and covariance matrices.
-Cluster means are placed with a minimum Wasserstein-2 distance constraint to ensure adequate separation between clusters and to prevent trivial overlap cases.
-The exact minimum distance value varies between 0.5 and 1.0 and is randomly selected for each component independently in order to promote data diversity.
+The goal is to create a continuous data type where each cluster follows a multivariate Gaussian distribution with carefully designed means and covariance matrices.
+This construction process is incremental.
+Starting with a mixture containing $k$ components, the addition of a new cluster involves initially placing it at position 0 and then shifting it in a randomly chosen direction.
 The covariance structure for each Gaussian is generated through eigendecomposition, with eigenvalues sampled from a predefined range of $[0.005,0.05]$ to control the shape and orientation of the clusters.
@@
 These constraints are activated with probabilities of 0.25 and 0.2, respectively.
-The process of constructing Gaussian mixture is incremental.
-Starting with a mixture containing $k$ components, the addition of a new cluster involves initially placing it at position 0 and then shifting it in a randomly chosen direction until the distance requirement is satisfied.
+To ensure adequate separation between clusters and to prevent trivial overlap, we apply a Wasserstein-2 distance constraint, requiring that the means $\mu_{i}$ and $\mu_{j}$ of any two clusters, along with their corresponding covariance matrices $\Sigma_{i}$ and $\Sigma_{j}$, are separated by at least a threshold $T$, as follows:
+    $\|\mu_{i}-\mu_{j}\|^{2}_{2}+tr(\Sigma_{i}+\Sigma_{j}-2(\Sigma_{i}^{\frac{1}{2}}\Sigma_{j}\Sigma_{i}^{\frac{1}{2}})^{\frac{1}{2}})\geq T$ (7)
+    The exact minimum distance value $T$ varies between 0.5 and 1.0 and is randomly selected for each component independently in order to promote data diversity.
 This entire approach allows us to model a variety of cluster geometries, from spherical to highly elongated elliptical shapes.
@@
     Categorical.
-We also incorporate categorical features alongside continuous ones by sampling from categorical distributions that are biased toward certain categories for each cluster.
-To generate these varied categorical probability patterns, we use the Dirichlet distribution.
+To enrich the generated data, we incorporate categorical features alongside continuous ones by sampling from categorical distributions that are biased toward certain categories for each cluster.
+To produce these varied categorical probability patterns, we use the Dirichlet distribution.
+The probability of including categorical features is controlled by a categorical_chancecategorical\_chance parameter, set to 0.3.
+Up to 3 categorical variables may be added, each having between 2 and max_categoriesmax\_categories possible values, defined as 5.
 The resulting categorical variables are then converted to one-hot encoding and combined with the continuous features, producing mixed-type datasets that more closely resemble real-world tabular data.
-The probability of including categorical features is controlled by a $categorical\_chance$ parameter, set to 0.3.
-Up to 3 categorical variables may be added, each having between 2 and $max\_categories$ possible values, defined as 5 .
   3. 3.
@@
 In addition, standardization is applied between residual layers to ensure more stable transformations.
-To help preserve cluster separability, we append one-hot vectors encoding the component identity of each data point to the numerical variables before passing them through the random neural network.
-After the transformation, these extra dimensions are removed using PCA, restoring the data to its original dimensionality.
+To help preserve cluster separability, we append one-hot vectors indicating the component identity to the numerical variables before passing them through the random neural network.
+After the transformation, these extra dimensions are removed using the PCA algorithm, restoring the data to its original dimensionality.
 This comprehensive approach enables the simulation of intricate, non-linear data manifold structures often presented in real-world clustering problems without introducing degenerate configurations.
-Afterwards, all features undergo standardization and scaling to ensure numerical stability during training.
-Continuous features are normalized to the range $[-1,1]$, while ensuring that the relative separation between clusters is preserved.
-By training on this diverse collection of synthetic datasets-ranging from well-separated Gaussian clusters to complex, transformed manifolds with mixed feature types-our model learns to identify meaningful cluster structures across a wide spectrum of data distributions.
-This comprehensive prior fitting enables ZEUS to adapt to previously unseen datasets at inference time without additional training.
+All features undergo standardization and scaling to ensure numerical stability during training.
+Continuous features are normalized to the range [1,1][-1,1], while ensuring that the relative separation between clusters is preserved.
+By training on this diverse collection of synthetic datasets-ranging from well-separated Gaussian clusters to complex, transformed manifolds with mixed feature types ZEUS learns to identify meaningful cluster structures across a wide spectrum of data distributions.
+This enables it to adapt to previously unseen datasets at inference time without additional training.
 Appendix C Pre-training plots
@@
 The first plot corresponds to the Gaussian-categorical datasets referred to as Synthetic Gaussian, while the second illustrates their NN transformed variants, named Synthetic Transformed.
-The data generation procedures are described in more detail in Appendix B.
+The data generation procedures are described in more detail in Appendix B.2.
 Both plots illustrate that the quality of the ZEUS representation improves over time, indicating that the model is learning new patterns, as evidenced by the increasing ARI during the pre-training process.
@@
 19 1493 11 0 11 0 7
-Tables 5, 6, and 7 provide detailed information about the datasets used in the experimental analysis.
+Tables 6, 7, and 8 provide detailed information about the datasets used in the experimental analysis.
 Each table includes the number of set instances, the number of categorical and numerical features, the total dimensionality, and the overall number of categories, represented by the length of the one-hot encoded vectors.
@@
   1. 1.
-    scikit-learn - used for k-means and GMM
+    scikit-learn - used for k-means and GMM,
   2. 2.
-    https://github.com/vlukiyanov/pt-dec - implementation of the DEC
+    https://github.com/vlukiyanov/pt-dec - implementation of the DEC,
   3. 3.
-    https://github.com/dawnranger/IDEC-pytorch - source code for the IDEC
+    https://github.com/dawnranger/IDEC-pytorch - source code for the IDEC,
   4. 4.
+    https://github.com/jsvir/idc - official implementation of the IDC method,
+  5. 5.
+  6. 6.
+    https://github.com/PriorLabs/tabpfn-extensions - a library that extends TabPFN functionality to a wide spectrum of machine learning tasks, including unsupervised ones,
+  7. 7.
+    https://github.com/clabrugere/pytorch-scarf - code repository for SCARF.
 To ensure fair comparison, hyperparameters are chosen to maximize the performance of each method with respect to their overall average rank across 5 random seeds.
 For this reason, all numerical features are preprocessed using a standard scaler prior to the training phase.
-Most parameters of the k-means and GMM methods are left at their default values. Only the $n\_init$ option was increased to 100 for k-means and 50 for GMM in order to improve stability.
+Most parameters of the k-means and GMM methods are left at their default values. Only the n_initn\_init option was increased to 100 for k-means and 50 for GMM in order to improve stability.
 As mentioned in Section 3.1, a standard autoencoder with hidden layers $[500,500,2000]$ is used for each AE-based method.
@@
 Changing other hyperparameters, including the learning rate, optimizer, and clustering loss weight $\gamma$, generally did not lead to improved results.
-Therefore, the remaining parameters were left at their default values as proposed by the authors of the respective repositories.
-Table 8: Real/OpenML datasets - ARI
-ID KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
-14 38.74 45.90 40.94 45.36 49.60 36.78 45.04 50.56
-15 82.85 71.38 61.44 71.74 86.29 87.48 24.38 81.28
-16 55.62 64.75 61.66 69.97 68.56 55.61 55.27 74.03
-18 54.86 46.88 51.45 51.47 50.00 53.54 49.93 51.63
-22 35.48 47.39 31.09 50.80 50.02 35.69 28.38 56.05
-35 70.68 70.68 70.22 70.48 68.63 64.58 56.57 85.12
-51 28.35 -2.54 37.82 34.13 38.13 34.58 35.13 43.66
-53 45.21 5.17 39.14 40.62 43.50 41.91 21.32 35.76
-56 57.79 58.49 57.10 58.31 56.97 55.20 55.66 66.41
-61 62.01 90.39 59.79 60.21 57.84 60.14 51.37 85.15
-187 89.75 93.09 82.87 84.86 85.02 85.10 55.67 88.19
-377 56.70 58.47 62.06 59.52 63.03 59.01 59.79 55.30
-458 95.08 96.97 64.33 75.70 94.58 63.29 68.31 99.19
-481 58.51 2.77 36.84 48.03 57.28 52.93 19.67 8.68
-694 35.91 43.30 35.26 36.56 26.30 36.02 30.76 33.38
-721 43.28 43.28 13.12 25.12 17.75 28.77 -0.03 43.28
-733 50.97 46.29 38.39 55.11 67.47 73.51 26.26 74.97
-745 57.63 55.44 78.35 75.52 46.37 50.94 57.81 73.85
-756 55.70 4.46 72.81 53.40 58.60 65.58 44.14 50.31
-796 49.62 2.26 76.54 52.19 50.07 55.92 43.40 13.99
-820 50.16 36.77 44.29 49.24 51.21 49.30 24.51 28.87
-840 39.05 48.37 53.60 46.72 49.14 39.16 27.95 27.66
-854 76.14 -0.21 62.33 60.83 35.05 70.52 72.25 76.14
-1462 1.32 0.31 10.03 10.17 1.00 3.80 8.48 92.28
-1495 96.81 98.40 72.74 82.80 96.81 91.28 77.51 7.35
-1499 77.33 62.99 73.68 75.28 69.99 78.48 29.60 82.40
-1510 67.07 78.02 63.73 59.77 71.32 69.84 69.47 74.26
-1523 21.16 41.21 25.48 24.57 28.16 23.05 18.50 19.60
-4153 55.78 58.03 59.15 46.23 56.41 58.44 39.43 62.29
-40496 54.02 34.79 42.35 37.51 49.74 42.35 37.33 32.25
-40682 58.32 86.29 57.89 59.22 57.39 86.05 40.58 53.15
-40705 43.49 37.92 0.73 30.63 47.90 46.28 10.71 45.45
-42261 62.01 90.39 60.13 58.15 59.78 56.78 49.09 85.15
-42585 60.84 30.52 51.26 60.71 91.79 43.40 38.30 95.05
-Mean 55.54 48.49 51.43 53.56 55.93 54.57 40.37 57.43
-Mean-rank 3.90 4.57 4.72 4.29 3.96 4.31 6.65 3.60
-Top-3 16 14 9 8 18 13 2 21
-Top-1 4 8 4 0 3 1 0 12
-Table 9: Synthetic gaussian datasets - ARI
-ID KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
-0 86.16 82.13 83.44 79.66 87.06 77.26 58.57 87.65
-1 88.66 87.36 81.13 79.61 85.84 75.62 54.10 72.04
-2 81.30 26.46 71.57 78.04 86.65 78.91 48.59 82.96
-3 95.02 36.77 92.03 76.44 92.74 89.81 62.67 95.76
-4 90.64 92.97 82.65 77.80 91.61 84.51 69.70 92.85
-5 89.83 25.85 86.52 69.06 90.01 88.31 67.70 90.72
-6 96.21 98.07 85.99 94.26 94.24 92.39 73.15 97.89
-7 92.59 30.52 81.73 79.06 91.35 89.80 58.10 93.60
-8 88.29 89.62 81.15 76.53 87.56 78.37 50.30 84.00
-9 93.36 98.54 68.58 85.42 91.14 85.29 56.89 94.55
-10 79.26 93.32 72.75 90.90 72.54 73.51 48.47 87.32
-11 93.60 98.48 90.77 89.11 95.03 92.19 64.24 97.45
-12 86.82 77.45 73.64 71.40 86.04 77.93 58.45 86.04
-13 96.77 98.07 94.05 90.15 96.11 85.66 78.20 96.17
-14 90.73 43.20 79.65 69.09 92.68 77.93 67.84 94.60
-15 93.58 96.23 85.74 89.30 94.43 89.41 72.88 96.68
-16 94.70 92.45 90.98 89.41 94.53 88.14 85.86 87.19
-17 85.26 87.10 62.23 79.52 84.08 75.00 42.55 86.29
-18 88.77 94.83 84.47 84.74 87.39 83.46 67.96 83.01
-19 86.54 89.11 76.18 78.46 86.06 67.99 70.50 73.85
-Mean 89.90 76.93 81.26 81.40 89.36 82.57 62.84 89.03
-Mean-Rank 2.60 3.35 5.40 5.45 3.10 5.50 7.70 2.90
-Top-3 16 13 0 1 15 0 0 15
-Top-1 3 10 0 0 1 0 0 6
-Table 10: Synthetic transformed datasets - ARI
-ID KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
-0 60.56 79.33 46.31 68.98 76.93 46.35 29.42 87.42
-1 93.30 96.55 76.57 83.99 91.29 70.91 73.09 89.40
-2 89.76 96.31 57.79 65.99 94.42 66.34 48.34 95.88
-3 70.16 50.38 56.23 76.06 83.06 60.09 54.54 80.98
-4 44.32 80.63 47.60 44.77 68.39 50.12 49.37 73.66
-5 91.49 96.51 74.48 89.11 86.46 77.84 66.47 81.63
-6 60.98 81.63 61.97 76.29 75.58 61.89 69.84 87.02
-7 76.80 83.98 71.37 77.69 86.56 77.49 56.51 89.57
-8 89.77 77.10 85.53 84.42 93.33 92.96 78.87 98.11
-9 84.40 98.48 82.64 93.03 97.14 81.02 65.62 97.20
-10 82.47 91.91 61.00 86.16 87.58 46.08 39.27 86.42
-11 90.77 99.32 81.41 91.18 94.78 88.12 70.29 93.19
-12 67.64 91.80 52.87 63.56 85.45 60.61 55.59 87.75
-13 85.99 50.80 45.48 35.34 29.04 12.91 32.89 91.60
-14 93.52 54.56 15.81 22.79 94.12 38.21 15.10 98.49
-15 36.62 30.93 37.71 43.70 42.78 40.65 29.70 28.96
-16 95.62 65.11 90.13 94.97 97.47 95.32 98.27 99.01
-17 60.54 95.74 56.24 91.71 78.72 59.11 13.29 96.44
-18 63.92 17.39 58.47 70.21 54.61 47.93 10.01 79.70
-19 62.12 79.06 49.36 65.77 81.07 51.19 26.84 84.13
-Mean 75.04 75.88 60.45 71.29 79.94 61.26 49.17 86.33
-Mean-Rank 4.45 3.20 6.20 4.25 3.00 5.70 7.05 2.15
-Top-3 5 14 0 6 15 2 1 17
-Top-1 0 8 0 1 1 0 0 10
+Therefore, the remaining parameters were left at their default values as proposed by the authors of the respective codebases.
+For TabPFN, we use the get_embeddings_per_column method from the TabPFNUnsupervisedModel class, provided in the TabPFN Extensions repository [^1] https://github.com/PriorLabs/tabpfn-extensions, to extract per-column representations.
+These embeddings are then averaged across columns to obtain the final representation for each data point.
+For SCARF, we follow the procedure outlined in the example.ipynb notebook from the original repository [^2] https://github.com/clabrugere/pytorch-scarf.
+In both cases, we cluster the learned representations using $k$-means, consistent with our approach in ZEUS.
 Appendix F Extended experimental results
@@
 F.1 How effective is ZEUS for clustering?
-Tables 8, 9, and 10 contain the complete results for individual datasets from the experiments discussed in Section 3.2 in Tables 1 and 2, as presented in the main text.
+Tables 9, 10, and 11 contain the complete results for individual datasets from the experiments discussed in Section 3.2 in Tables 1 and 2, as presented in the main text.
 The reported values represent averages of ARI over 5 random seeds.
@@
 One noteworthy point is that ZEUS achieves strong performance in terms of the Top-3 and Top-1 statistics, being outperformed in this regard only in Table 10, where k-means and GMM score better.
-Another interesting observation is that GMM struggles with categorical features, which significantly worsens its average scores in Tables 9 and 10.
+Another interesting observation is that GMM struggles with categorical features, which significantly worsens its average scores in Tables 10 and 11.
+Table 9: Comparison of clustering quality using ARI metric on OpenML datasets (higher score indicate better performance).
+ID KM GMM AE-KM AE-GMM DEC IDEC IDC G-CEALS TabPFN SCARF ZEUS
+14 38.74 45.90 40.94 45.36 49.60 36.78 45.03 45.04 22.02 1.90 50.56
+15 82.85 71.38 61.44 71.74 86.29 87.48 83.20 24.38 74.31 75.01 81.28
+16 55.62 64.75 61.66 69.97 68.56 55.61 65.84 55.27 14.69 3.74 74.03
+18 54.86 46.88 51.45 51.47 50.00 53.54 44.40 49.93 51.50 5.02 51.63
+22 35.48 47.39 31.09 50.80 50.02 35.69 46.03 28.38 11.09 2.63 56.05
+35 70.68 70.68 70.22 70.48 68.63 64.58 77.99 56.57 52.24 74.61 85.12
+51 28.35 -2.54 37.82 34.13 38.13 34.58 38.93 35.13 38.26 -1.27 43.66
+53 45.21 5.17 39.14 40.62 43.50 41.91 33.03 21.32 26.61 24.24 35.76
+56 57.79 58.49 57.10 58.31 56.97 55.20 59.25 55.66 58.42 61.26 66.41
+61 62.01 90.39 59.79 60.21 57.84 60.14 58.30 51.37 77.20 51.19 85.15
+187 89.75 93.09 82.87 84.86 85.02 85.10 80.04 55.67 37.55 40.05 88.19
+377 56.70 58.47 62.06 59.52 63.03 59.01 62.39 59.79 66.89 14.45 55.30
+458 95.08 96.97 64.33 75.70 94.58 63.29 86.77 68.31 61.41 2.87 99.19
+481 58.51 2.77 36.84 48.03 57.28 52.93 43.17 19.67 2.77 1.79 8.68
+694 35.91 43.30 35.26 36.56 26.30 36.02 33.47 30.76 27.68 14.09 33.38
+721 43.28 43.28 13.12 25.12 17.75 28.77 12.74 -0.03 23.63 6.35 43.28
+733 50.97 46.29 38.39 55.11 67.47 73.51 68.88 26.26 -8.35 49.41 74.97
+745 57.63 55.44 78.35 75.52 46.37 50.94 59.29 57.81 1.26 16.75 73.85
+756 55.70 4.46 72.81 53.40 58.60 65.58 53.32 44.14 10.60 64.82 50.31
+796 49.62 2.26 76.54 52.19 50.07 55.92 79.59 43.40 0.88 58.25 13.99
+820 50.16 36.77 44.29 49.24 51.21 49.30 44.59 24.51 50.26 13.04 28.87
+840 39.05 48.37 53.60 46.72 49.14 39.16 19.59 27.95 8.89 32.13 27.66
+854 76.14 -0.21 62.33 60.83 35.05 70.52 64.81 72.25 -0.54 6.48 76.14
+1462 1.32 0.31 10.03 10.17 1.00 3.80 8.83 8.48 3.31 -0.03 92.28
+1495 96.81 98.40 72.74 82.80 96.81 91.28 83.96 77.51 47.11 5.10 7.35
+1499 77.33 62.99 73.68 75.28 69.99 78.48 67.82 29.60 5.82 24.93 82.40
+1510 67.07 78.02 63.73 59.77 71.32 69.84 70.55 69.47 54.59 63.57 74.26
+1523 21.16 41.21 25.48 24.57 28.16 23.05 31.35 18.50 2.45 21.04 19.60
+4153 55.78 58.03 59.15 46.23 56.41 58.44 50.44 39.43 27.80 36.21 62.29
+40496 54.02 34.79 42.35 37.51 49.74 42.35 27.41 37.33 31.80 39.78 32.25
+40682 58.32 86.29 57.89 59.22 57.39 86.05 17.95 40.58 86.94 17.86 53.15
+40705 43.49 37.92 0.73 30.63 47.90 46.28 31.38 10.71 -4.40 32.90 45.45
+42261 62.01 90.39 60.13 58.15 59.78 56.78 55.10 49.09 77.20 52.53 85.15
+42585 60.84 30.52 51.26 60.71 91.79 43.40 72.10 38.30 23.04 3.50 95.05
+Mean 55.54 48.49 51.43 53.56 55.93 54.57 52.28 40.37 31.32 26.95 57.43
+Mean-Rank 4.69 5.65 5.72 5.24 4.69 5.01 5.62 8.18 8.22 8.85 4.13
+Top-3 11 12 6 5 15 11 9 1 6 4 21
+Top-1 4 7 3 0 2 1 1 0 2 0 12
+Table 10: Evaluation of clustering quality with the ARI metric on Synthetic Gaussian datasets (higher scores reflect better performance).
+ID KM GMM AE-KM AE-GMM DEC IDEC IDC G-CEALS TabPFN SCARF ZEUS
+0 86.16 82.13 83.44 79.66 87.06 77.26 60.64 58.57 65.42 4.06 87.65
+1 88.66 87.36 81.13 79.61 85.84 75.62 44.44 54.10 40.26 4.21 72.04
+2 81.30 26.46 71.57 78.04 86.65 78.91 53.53 48.59 20.42 11.44 82.96
+3 95.02 36.77 92.03 76.44 92.74 89.81 65.03 62.67 50.57 22.29 95.76
+4 90.64 92.97 82.65 77.80 91.61 84.51 77.56 69.70 61.65 8.80 92.85
+5 89.83 25.85 86.52 69.06 90.01 88.31 70.38 67.70 20.51 10.31 90.72
+6 96.21 98.07 85.99 94.26 94.24 92.39 80.18 73.15 86.50 6.51 97.89
+7 92.59 30.52 81.73 79.06 91.35 89.80 46.44 58.10 26.62 10.14 93.60
+8 88.29 89.62 81.15 76.53 87.56 78.37 72.72 50.30 68.54 13.06 84.00
+9 93.36 98.54 68.58 85.42 91.14 85.29 76.64 56.89 65.02 11.37 94.55
+10 79.26 93.32 72.75 90.90 72.54 73.51 77.64 48.47 89.15 -0.04 87.32
+11 93.60 98.48 90.77 89.11 95.03 92.19 74.56 64.24 68.28 2.11 97.45
+12 86.82 77.45 73.64 71.40 86.04 77.93 60.55 58.45 55.27 19.11 86.04
+13 96.77 98.07 94.05 90.15 96.11 85.66 86.14 78.20 80.24 5.29 96.17
+14 90.73 43.20 79.65 69.09 92.68 77.93 72.80 67.84 12.93 8.16 94.60
+15 93.58 96.23 85.74 89.30 94.43 89.41 72.54 72.88 91.02 3.24 96.68
+16 94.70 92.45 90.98 89.41 94.53 88.14 77.56 85.86 59.58 4.85 87.19
+17 85.26 87.10 62.23 79.52 84.08 75.00 50.45 42.55 50.06 7.69 86.29
+18 88.77 94.83 84.47 84.74 87.39 83.46 61.36 67.96 74.90 7.69 83.01
+19 86.54 89.11 76.18 78.46 86.06 67.99 47.46 70.50 32.39 6.02 73.85
+Mean 89.90 76.93 81.26 81.40 89.35 82.57 66.43 62.84 55.97 8.32 89.03
+Mean-Rank 2.65 3.65 5.65 5.60 3.23 5.70 7.95 8.90 8.75 11.00 2.92
+Top-3 16 13 0 1 15 0 0 0 1 0 14
+Top-1 3 10 0 0 1 0 0 0 0 0 6
+Table 11: Assessment of clustering quality based on the ARI metric on Synthetic Transformed datasets (higher score indicate better performance).
+ID KM GMM AE-KM AE-GMM DEC IDEC IDC G-CEALS TabPFN SCARF ZEUS
+0 60.56 79.33 46.31 68.98 76.93 46.35 56.67 29.42 6.36 0.94 87.42
+1 93.30 96.55 76.57 83.99 91.29 70.91 66.59 73.09 16.59 1.22 89.40
+2 89.76 96.31 57.79 65.99 94.42 66.34 72.25 48.34 10.00 1.10 95.88
+3 70.16 50.38 56.23 76.06 83.06 60.09 65.13 54.54 15.36 1.99 80.98
+4 44.32 80.63 47.60 44.77 68.39 50.12 36.00 49.37 19.14 1.48 73.66
+5 91.49 96.51 74.48 89.11 86.46 77.84 60.00 66.47 15.12 1.25 81.63
+6 60.98 81.63 61.97 76.29 75.58 61.89 69.43 69.84 18.83 2.11 87.02
+7 76.80 83.98 71.37 77.69 86.56 77.49 76.75 56.51 18.94 3.20 89.57
+8 89.77 77.10 85.53 84.42 93.33 92.96 90.69 78.87 24.26 5.91 98.11
+9 84.40 98.48 82.64 93.03 97.14 81.02 91.02 65.62 9.01 1.66 97.20
+10 82.47 91.91 61.00 86.16 87.58 46.08 70.42 39.27 14.11 0.46 86.42
+11 90.77 99.32 81.41 91.18 94.78 88.12 80.75 70.29 18.31 1.72 93.19
+12 67.64 91.80 52.87 63.56 85.45 60.61 73.66 55.59 28.50 3.21 87.75
+13 85.99 50.80 45.48 35.34 29.04 12.91 30.22 32.89 35.78 6.01 91.60
+14 93.52 54.56 15.81 22.79 94.12 38.21 68.54 15.10 15.08 5.54 98.49
+15 36.62 30.93 37.71 43.70 42.78 40.65 32.64 29.70 16.68 0.63 28.96
+16 95.62 65.11 90.13 94.97 97.47 95.32 84.79 98.27 17.27 9.58 99.01
+17 60.54 95.74 56.24 91.71 78.72 59.11 82.54 13.29 7.33 -0.10 96.44
+18 63.92 17.39 58.47 70.21 54.61 47.93 64.15 10.01 3.79 0.65 79.70
+19 62.12 79.06 49.36 65.77 81.07 51.19 63.41 26.84 2.73 1.06 84.13
+Mean 75.04 75.88 60.45 71.29 79.94 61.26 66.78 49.17 15.66 2.48 86.33
+Mean-Rank 4.80 3.50 6.85 4.50 3.20 6.35 6.05 7.80 9.75 11.00 2.20
+Top-3 4 14 0 6 15 2 1 1 0 0 17
+Top-1 0 8 0 1 1 0 0 0 0 0 10
 F.2 Are ZEUS’s assignments well calibrated?
-Table 11, 12 and 13 contain detailed extension of Table 3
-Table 11: Real/OpenML datasets - Brier
+Tables 12, 13 and 14 contain detailed extension of Table 3 from the main part of the paper.
+Similar to the Tables in Appendix F.1, these also include additional rows: Mean, Mean-Rank, Top-3, and Top-1, which aggregate the results presented in each table.
+ZEUS consistently attains at least the second position across all statistics in every table presented in this section.
+For the Synthetic Transformed datasets, it is undeniably the best.
+However, for the OpenML datasets, it is outperformed by IDEC, and for the Synthetic Gaussian datasets, GMM takes the lead in the Top-1 metric, while for the other statistics, k-means performs better.
+Table 12: Soft clustering performance of ZEUS versus competing methods on real-world datasets, measured by Brier score (lower is better).
 ID KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
@@
 Top-1 3 5 1 1 6 8 1 8
-Table 12: Synthetic gaussian datasets - Brier
+Table 13: Soft clustering performance (Brier score) of ZEUS compared to baseline methods on Synthetic Gaussian datasets (lower score reflect better quality).
 ID KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
@@
 Top-1 2 11 0 1 0 0 0 6
-Table 13: Synthetic transformed datasets - Brier
+Table 14: Evaluation of soft clustering quality on Synthetic Transformed datasets using Brier score (lower score indicate better performance).
 ID KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
@@
 F.3 How helpful is regularization for ZEUS?
-Tables 14, 15 and 16 provide extended versions of Table 4 , which analyzes the impact of different combinations of regularization functions.
-Table 14: Real/OpenML datasets - ablations ARI
+Tables 15, 16 and 17 provide extended versions of Table 4 , which analyzes the impact of different combinations of regularization functions.
+Their structure is analogous to the other tables presented in Appendix F.
+The conclusions that can be drawn from these extended tables, along with their statistical summaries, are consistent with our original claims.
+The model with both $\mathcal{L}_{sep}$ and $\mathcal{L}_{cp}$ regularizers performs best on real-world data collections and consistently ranks at least second in the considered statistics for synthetic data.
+On these generated datasets, it is frequently outperformed by the variant that includes the $\mathcal{L}_{cp}$ component alone.
+By contrast, the remaining two approaches clearly grapple with proper clustering of the synthetic datasets.
+Table 15: Impact of regularization components on OpenML datasets, evaluated using the ARI score (higher is better).
 ID $\mathcal{L}_{prob}$ $\mathcal{L}_{prob}+\mathcal{L}_{sep}$ $\mathcal{L}_{prob}+\mathcal{L}_{cp}$ $\mathcal{L}_{prob}+\mathcal{L}_{sep}+\mathcal{L}_{cp}$
@@
 Top-1 2 9 6 15
-Table 15: Synthetic gaussian datasets - ablations ARI
+Table 16: Effect of regularization components on Synthetic Gaussian datasets, assessed by the ARI metric (higher score indicate better quality).
 ID $\mathcal{L}_{prob}$ $\mathcal{L}_{prob}+\mathcal{L}_{sep}$ $\mathcal{L}_{prob}+\mathcal{L}_{cp}$ $\mathcal{L}_{prob}+\mathcal{L}_{sep}+\mathcal{L}_{cp}$
@@
 Top-1 3 3 7 7
-Table 16: Synthetic transformed datasets - ablations ARI
+Table 17: Regularisation impact on Synthetic Transformed datasets, measured by ARI (higher is better).
 ID $\mathcal{L}_{prob}$ $\mathcal{L}_{prob}+\mathcal{L}_{sep}$ $\mathcal{L}_{prob}+\mathcal{L}_{cp}$ $\mathcal{L}_{prob}+\mathcal{L}_{sep}+\mathcal{L}_{cp}$
@@
 Top-1 2 5 6 7
-Appendix G Licensing and Third-Party Assets.
+F.4 What data-generating prior is optimal?
+Tables 18, 19, and 20 expand upon the findings presented in Table 5 by examining the influence of various data-generating priors across all datasets within each group.
+Their format is consistent with other tables in Appendix F.
+Table 18: Evaluation of clustering quality with the ARI metric on real-world datasets for different combinations of data-generating probabilistic models (higher scores reflect better performance).
+ID Gauss. + Cat. NN-transf. + Cat. Gauss. + NN-transf. Gauss. + NN-transf. + Cat.
+14 36.81 49.84 49.84 50.56
+15 85.52 85.57 83.46 81.28
+16 57.15 73.05 76.01 74.03
+18 42.83 46.48 55.29 51.63
+22 46.45 62.23 60.63 56.02
+35 87.48 76.48 84.25 85.12
+51 -0.52 38.26 39.76 43.66
+53 37.57 29.79 27.38 35.76
+56 48.40 56.35 59.92 66.41
+61 68.44 62.26 85.08 85.15
+187 89.77 94.87 88.22 88.19
+377 35.12 58.34 58.95 55.30
+458 47.43 98.79 98.47 99.19
+481 6.46 6.94 3.58 8.68
+694 27.75 19.90 27.69 33.38
+721 43.28 43.28 38.13 43.28
+733 73.48 -7.69 75.77 74.97
+745 4.46 67.51 28.11 73.85
+756 4.46 59.53 4.46 50.31
+796 14.80 5.94 15.29 13.99
+820 19.26 36.77 33.71 28.87
+840 -0.83 0.69 2.57 27.66
+854 24.32 73.92 12.56 76.14
+1462 1.92 92.56 92.56 92.28
+1495 -0.29 0.03 55.17 7.35
+1499 58.62 74.43 71.07 82.40
+1510 72.50 72.95 70.08 74.26
+1523 20.38 18.69 22.01 19.60
+4153 41.49 46.33 56.13 62.29
+40496 33.20 31.69 30.61 32.25
+40682 64.49 58.59 77.21 53.15
+40705 29.52 39.74 37.15 45.45
+42261 66.34 62.26 81.76 85.15
+42585 92.06 94.29 64.95 95.05
+Mean 40.59 50.90 52.00 57.43
+Mean-Rank 3.10 2.56 2.46 1.88
+Top-3 17 27 26 31
+Top-1 3 5 8 16
+Table 19: Effect of prior selection on ZEUS’s ARI performance on Synthetic Gaussian datasets (higher scores indicate better performance).
+ID Gauss. + Cat. NN-transf. + Cat. Gauss. + NN-transf. Gauss. + NN-transf. + Cat.
+0 88.78 87.09 82.89 87.65
+1 87.74 79.51 80.38 72.04
+2 91.30 87.30 20.50 82.96
+3 95.08 93.27 44.56 95.76
+4 94.40 94.46 92.48 92.85
+5 91.43 90.08 30.45 90.72
+6 98.05 97.22 98.02 97.89
+7 94.56 94.43 26.90 93.60
+8 87.65 87.95 88.05 84.00
+9 95.92 92.70 94.79 94.55
+10 90.99 86.78 90.07 87.32
+11 96.86 96.52 96.74 97.45
+12 88.21 86.25 85.55 86.04
+13 96.07 93.39 93.81 96.17
+14 95.42 92.12 32.24 94.60
+15 97.71 95.84 97.14 96.68
+16 95.24 94.49 95.50 87.19
+17 89.13 84.54 86.66 86.29
+18 90.14 87.66 89.37 83.01
+19 87.43 76.30 79.00 73.85
+Mean 92.61 89.90 75.25 89.03
+Mean-Rank 1.35 3.05 2.80 2.80
+Top-3 20 13 12 15
+Top-1 14 1 2 3
+Table 20: Influence of the chosen data-generating prior on ZEUS performance over Synthetic Transformed datasets, reported in terms of ARI (larger values denote better clustering quality).
+ID Gauss. + Cat. NN-transf. + Cat. Gauss. + NN-transf. Gauss. + NN-transf. + Cat.
+0 61.54 87.33 88.33 87.42
+1 84.60 95.73 94.60 89.40
+2 83.73 93.89 97.59 95.88
+3 45.52 75.48 52.79 80.98
+4 76.52 73.72 81.24 73.66
+5 72.95 83.07 79.33 81.63
+6 80.87 83.20 83.19 87.02
+7 88.97 95.91 91.18 89.57
+8 97.55 98.09 36.45 98.11
+9 94.25 96.86 98.29 97.20
+10 40.64 81.56 88.46 86.42
+11 93.42 98.57 98.98 93.19
+12 83.00 90.91 92.21 87.75
+13 80.76 90.87 94.12 91.60
+14 98.19 98.49 13.51 98.49
+15 28.32 28.47 29.06 28.96
+16 98.40 99.05 18.65 99.01
+17 87.75 98.21 98.21 96.44
+18 5.50 83.07 3.79 79.70
+19 64.25 88.36 85.72 84.13
+Mean 73.34 87.04 71.29 86.33
+Mean-Rank 3.65 2.00 2.08 2.27
+Top-3 6 20 16 18
+Top-1 0 6 9 3
+Appendix G Additional experiments
+In this section, we present additional ablation studies and experiments related to ZEUS, including analyses regarding its robustness, architectural choices, clustering strategies and performance in special-case scenarios.
+G.1 Robustness to noise and outliers
+In order to assess the robustness of our model under noise and outlier conditions, we consider two alternative evaluation setups:
+  • ZEUS + noise, where we add Gaussian noise with a standard deviation of 0.05 to each data point before the ZEUS forward pass,
+[…]
+Table 21 displays the outcomes of these experiments, alongside the performance of the vanilla ZEUS model.
+Table 21: Impact of Gaussian noise and anomalies on clustering performance.
+Metric Dataset group ZEUS+noise ZEUS+anomalies ZEUS
+ARI Real 55.88 55.83 57.43
+ARI Syn. Gauss. 88.01 88.21 89.03
+ARI Syn. Transf. 84.82 85.17 86.33
+Rank Real 2.24 1.93 1.84
+Rank Syn. Gauss. 2.45 2.10 1.45
+Rank Syn. Transf. 2.45 2.30 1.25
+Overall, the average performance in the perturbed setups is not significantly lower than that of the vanilla ZEUS model, highlighting a degree of stability of our method in the presence of noise and outliers.
+G.2 Impact of token dimension
+The choice of token dimension also has a substantial influence on the performance and generalizability of ZEUS.
+This is confirmed by an experiment in which we evaluated dimensions of 32, 128, 512, and 768 while keeping the standard setup unchanged.
+Table 22 reports the clustering performance of these variants, denoted as ZEUS-32, ZEUS-128, ZEUS-512, and ZEUS-768, respectively.
+Table 22: Assessment of different token dimensions in the ZEUS architecture.
+Metric Dataset group ZEUS-32 ZEUS-128 ZEUS-512 ZEUS-768
+ARI Real 44.10 47.81 57.43 50.79
+ARI Syn. Gauss. 69.50 83.04 89.03 91.92
+ARI Syn. Transf. 67.57 81.62 86.33 87.21
+Rank Real 2.90 2.96 1.90 2.25
+Rank Syn. Gauss. 4.00 3.00 1.78 1.23
+Rank Syn. Transf. 3.90 2.78 1.70 1.63
+As one can infer, increasing the dimensionality leads to a consistent improvement in performance for synthetic data, whereas for OpenML datasets, there is a drop between dimensions 512 and 768, which may be caused by overfitting to the prior.
+G.3 Effect of the output clustering method
+Since ZEUS is a method for learning representations suitable for clustering, in principle, any clustering algorithm could be used to separate the resulting clusters.
+However, our approach has the additional feature of being inherently designed to form spherical clusters; therefore, $K$-Means (ZEUS-KM) should, in theory, already be sufficient.
+The primary advantage of $K$-means lies in its speed, which is a crucial factor for us during inference.
+Moreover, as demonstrated in the Table 23, its clustering quality (measured by ARI) is comparable to the more computationally intensive Gaussian Mixture Model (ZEUS-GMM).
+For reference, we also include results from a simplified variant of GMM with fixed identity covariance matrices (ZEUS-SGMM), as employed in Section 3.3.
+All experiments are conducted using the same checkpoint and datasets as in Section 3.2.
+Table 23: Evaluation of clustering methods applied to the learned ZEUS representations.
+Metric Dataset group ZEUS-KM ZEUS-SGMM ZEUS-GMM
+ARI Real 57.43 57.33 56.45
+ARI Syn. Gauss. 89.03 88.18 89.46
+ARI Syn. Transf. 86.33 85.57 87.22
+Rank Real 1.94 2.04 2.01
+Rank Syn. Gauss. 1.90 2.33 1.78
+Rank Syn. Transf. 2.13 2.35 1.53
+Analysis of the average ARI scores indicates that ZEUS’s output is well-suited to the evaluated approaches, as there are no significant differences in their results.
+Deep clustering techniques such as DEC or IDEC could also potentially be employed.
+Nevertheless, their reliance on intensive, dataset-specific training contradicts ZEUS’s objective of delivering efficient, zero-shot representations.
+G.4 Integrating TabICL transformer into ZEUS
+To mitigate one of the main limitations of our approach, namely the restricted number of input features and samples, we investigate whether it is feasible to replace the original TabPFN transformer architecture with a more recent design, such as the one used in TabICL [18].
+Specifically, we adopt the TabICL transformer with its default hyperparameters and integrate it into our unsupervised pipeline (ZEUS + TabICL) in exactly the same manner as described in Section 3.1.
+The resulting clustering quality, alongside a comparison with the original ZEUS model, are presented in Table 24.
+Table 24: Performance of ZEUS vs. its variant with the TabICL transformer backbone.
+Metric Dataset group ZEUS ZEUS + TabICL
+ARI Real 57.43 52.62
+ARI Syn. Gauss. 89.03 89.81
+ARI Syn. Transf. 86.33 85.97
+Rank Real 1.38 1.62
+Rank Syn. Gauss. 1.60 1.40
+Rank Syn. Transf. 1.75 1.25
+Overall, ZEUS+TabICL performs competitively with the original ZEUS across all evaluated data collections and even surpasses it in rank on synthetic data.
+Nonetheless, its average ARI on real-world datasets is noticeably lower.
+Future work should investigate whether this is due to random seed variability, our incomplete understanding of the TabICL architecture, potential overfitting to the prior, or the need for a more expressive prior.
+G.5 Special-case scenarios
+Finally, we examine the behavior of our model in several extreme scenarios. To this end, we construct four synthetic two-dimensional datasets designed to highlight specific challenges:
+  • Cluster-Imbalance: Three clusters with 1000, 300, and 50 samples, respectively.
+  • Covariance-Scaled: Three equally sized clusters (500 samples each) with covariance matrices scaled by factors of 1, 6, and 36.
+  • Moons: Generated using make_moons from scikit-learn with 1000 samples and 0.1 noise.
+  • Connection: Two spherical Gaussians linked by a degenerate linear Gaussian.
+We report a comparison of ZEUS performance against $K$-Means and GMM in Table 25.
+Table 25: Comparison of clustering quality on synthetic special-case datasets, assessed by the ARI metric (higher scores indicate better quality).
+KM GMM ZEUS
+Cluster-Imbalance 50.12 96.75 96.80
+Covariance-scaled 79.68 96.45 94.72
+Moons 48.67 49.23 98.80
+Connection 72.76 89.30 83.16
+In summary, ZEUS outperforms $K$-means across all cases and performs on par with GMM.
+Notably, on the Moons dataset, ZEUS is the only method that successfully separates the clusters.
+This suggests that these synthetic examples do not represent genuine failure cases for our approach.
+A more thorough analysis of datasets where ZEUS underperforms remains an important direction for future work.
+Appendix H Licensing and Third-Party Assets.
 Our method builds upon the publicly available TabPFN codebase, which is released under the Apache 2.0 License.
@@
 Although ZEUS is designed as a general-purpose tool for clustering tabular data, its use in high-stakes domains like healthcare, finance, or criminal justice carries inherent risks.
-In such contexts, automated grouping of individuals—especially without labels or fairness constraints—can lead to biased or opaque outcomes.
+In such contexts, automated grouping of individuals - especially without labels or fairness constraints - can lead to biased or opaque outcomes.
 We therefore recommend that any deployment of ZEUS in sensitive applications be accompanied by fairness assessment and expert review.